### v1

# USTC-ICTS/PCFT-25-37

Abstract
We introduce a novel, systematic method for the complete symbolic reduction of multi-loop Feynman integrals, leveraging the power of generating functions. The differential equations governing these generating functions naturally yield symbolic recurrence relations. We develop an efficient algorithm that utilizes these recurrences to reduce integrals to a minimal set of master integrals. This approach circumvents the exponential growth of traditional Integration-by-Parts relations, enabling the reduction of high-rank, multi-loop integrals critical for state-of-the-art calculations in perturbative quantum field theory.
I Introduction
Loop-level Feynman diagrams play a central role in perturbative quantum field theories (QFTs). In the era of precision physics, The current era of precision physics presents an urgent demand for the calculation of multi-loop amplitudes, which are crucial for collider phenomenology, gravitational wave predictions from black hole scattering, as well as formal aspects of QFT. In multi-loop computations, a key step is to reduce all Feynman integrals to a finite, linearly independent set consisting of master integrals. This step, known as Feynman integral reduction, is frequently the computationally intensive bottleneck of perturbative QFT calculation. Conventionally, integral reduction is carried out using the integration-by-parts (IBP) identities [1, 2] via the linear reduction algorithm named after Laporta [3]. However, for most cutting-edge multi-loop computations, the IBP method inevitably involves a vast linear system, which is prohibitively difficult to solve via linear reduction.
There are several public software packages for Feynman integral reduction, implementing the Laporta algorithm along with various modern techniques: AIR [4], FIRE [5, 6, 7, 8], LiteRed [9, 10], Reduze [11, 12], FiniteFlow [13], Kira [14, 15, 16, 17, 18], FireFly [19, 20], RaTracer [21], Blade [22, 23, 24], and NeatIBP [25, 26]. Notable advances that have boosted linear reduction include the finite-field method [27], the application of computational algebraic geometry [28, 29, 25, 26], and the block-triangular form [22, 23, 24]. An alternative framework is provided by intersection theory [30, 31, 32], formulating Feynman integral reduction in terms of twisted cohomology. Despite these developments, the capabilities of existing reduction methods are still insufficient to meet the demands of upcoming precision experiments, such as the High-Luminosity Large Hadron Collider (HL-LHC), the Laser Interferometer Space Antenna (LISA), the Taiji and Tianqin projects.
The main bottleneck of integration-by-parts (IBP) reduction arises in multi-loop integrals with high-degree numerators or high-power propagators, where the number of IBP relations grows exponentially. This issue becomes particularly severe in recent cutting-edge multi-loop computations, such as the $5$ pm $1$ SF black-hole scattering [33] and the three-loop five-point Feynman integrals [34]. Tackling reduction problems at the three- and four-loop orders, or for two-loop multi-leg processes, necessitates new conceptual breakthroughs.
Recurrence rules [35, 9] provide an efficient alternative to IBP reductions, particularly for the challenging task of reducing multi-loop integrals with high-rank numerators or elevated propagator powers. The implements of recurrence rules include methods based on gröbner basis [36, 37, 35, 38, 39] and heuristic algorithms [9, 10]. However, a significant drawback was the computational cost associated with the recurrence rule derivation, which has limited their widespread adoption.
In this work, we present a novel and highly efficient method, based on generating functions [40, 41], to systematically derive complete symbolic recurrence relations for the reduction of Feynman integrals. By expressing Feynman integrals in terms of generating functions, integration-by-parts (IBP) relations are translated into differential equations for these generating functions. The reduction of Feynman integrals is then transformed into the task of reducing a differential operator with respect to these differential equations. We introduce a new approach for differential operator reduction, based on operator grading and integrability conditions, which allows us to efficiently tobtain symbolic reduction rules for arbitrary Feynman integrals.
This paper is organized as follows. We first review the definition of generating functions for Feynman integral reduction. We then focus on our new systematic method for obtaining symbolic reduction rules. As illustrations, we apply this method to three multi-loop reduction examples: the sunset diagram, and planar and non-planar double-box diagrams.
During the preparation of this manuscript, we became aware of the recent paper [42], which presents an efficient method for generating Feynman integral reduction rules using syzygies. In this work, we emphasize that our method uses only fundamental IBP relations and does not rely on syzygies.
II Generating Functions for Feynman integrals
For $L$-loop integrals with $E$ independent external momenta, a family of scalar FIs can be represented as
$I(\vec{\nu})=\int\prod_{i=1}^{L}\frac{\mathrm{d}^{D}\ell_{i}}{\mathrm{i}\pi^{D/2}}\frac{\mathcal{D}_{K+1}^{-\nu_{K+1}}\cdots\mathcal{D}_{N}^{-\nu_{N}}}{\mathcal{D}_{1}^{\nu_{1}}\cdots\mathcal{D}_{K}^{\nu_{K}}},\penalty 10000\ \penalty 10000\ \penalty 10000\$(1)
where $D=4-2\epsilon$ represents the spacetime dimension, $\ell_{i}$ are the loop momenta, $\mathcal{D}_{1},\ldots,\mathcal{D}_{K}$ are the inverse propagators, $\mathcal{D}_{K+1},\ldots,\mathcal{D}_{N}$ are the irreducible scalar products (ISPs) introduced to form a complete basis and $N={L(L+1)\over 2}+LE$. In (1), $\nu_{1},\ldots,\nu_{K}$ can be arbitrary integers (i.e, positive, negative or zero), while $\nu_{K+1},\ldots,\nu_{N}$ can only be non-positive integers. An integral is said to belong to the top-sector of this family if all its propagator indices are positive, i.e., $\nu_{1},\ldots,\nu_{K}>0$.
To organize all integrals within a sector, we introduce a generating function defined as:
$G_{\vec{\mu}}(\eta)=\int\prod_{i=1}^{L}\frac{\mathrm{d}^{D}\ell_{i}}{\mathrm{i}\pi^{D/2}}{e^{\sum_{j=1}^{N}(1-\mu_{j})\eta_{j}s_{0}^{-1}\mathcal{D}_{j}}\over\prod_{i=1}^{N}(\mathcal{D}_{i}-s_{0}\eta_{i})^{\mu_{i}}},\penalty 10000\ \penalty 10000\ \penalty 10000\$(2)
where $\vec{\mu}=(\mu_{1},...,\mu_{N})$ consists of $\mu_{i}=0,1$ and $s_{0}$ is an auxiliary scale that make $\eta_{i}$ dimensionless. For top-sector, we set $\mu_{1},\cdots,\mu_{K}=1$ and $\mu_{K+1},\cdots,\mu_{N}=0$ and $G_{\vec{\mu}}$ can be expanded as:
$G_{\vec{\mu}}(\vec{\eta})$ $=$ $\sum_{\vec{n}\geq 0}\vec{\eta}^{\vec{n}}F_{\vec{n}},\penalty 10000\ \penalty 10000\ \penalty 10000\$(3)
with
$F_{\vec{n}}$ $=$ $\left(\prod_{j=K+1}^{N}{1\over n_{j}!}\right)s_{0}^{\sum_{i=1}^{K}n_{i}-\sum_{j=K+1}^{N}n_{j}}$ (4)
$\times$ $I(...,n_{K}+1,-n_{K+1},...,-n_{N})\penalty 10000\ .$
The formalism (3) has mapped Feynman integrals to the expansion coefficients of generating function, labeled by the lattice points $\vec{n}$. We define the degree of $F$ as $\sum_{i=1}^{N}n_{i}$. Consequently, the reduction problem of Feynman integrals is transformed into finding relations among these coefficients, which can be derived from differential equations(DEs) satisfied by the generating function.
Since the generating function encodes all Feynman integrals in the family, it is sufficient to construct IBP relations for the generating function alone, which we call as the fundamental IBP relations. These relations yield DEs of the form:
$\sum_{t}c_{t}\hat{O}_{t}G_{\vec{\mu}}(\vec{\eta})=B\penalty 10000\ ,\penalty 10000\ \penalty 10000\ \penalty 10000\ \hat{O}_{t}=\prod_{i=1}^{N}\eta_{i}^{(\vec{a}_{t})_{i}}{\partial^{(\vec{b}_{t})_{i}}\over\partial\eta_{i}^{(\vec{b}_{t})_{i}}\penalty 10000\ },\penalty 10000\$(5)
where $B$ is the function of generating functions of subsectors and $c_{t}$ are coefficients. We define the degree of an operator as $\sum_{i=1}^{N}(\vec{o})_{i}$, where the index vector is $\vec{o}\equiv\vec{b}-\vec{a}$. By substituting the expansion (3) to equation (5) and extracting the coefficient of $\vec{\eta}^{\vec{n}}$, we obtain
$\sum_{t}c_{t}\left(\prod_{j=1}^{N}\prod_{p=1}^{(\vec{b}_{t})_{j}}(n_{j}-(\vec{a})_{j}+p)\right)F_{\vec{n}+\vec{b}_{t}-\vec{a}_{t}}=\widetilde{B}$$penalty 10000\ \penalty 10000\ \penalty 10000\$(6)
where for simplicity, we have made the convention that $F_{\vec{n}}=0$ if any component $(\vec{n})_{i}<0$. We call equation (6) as the recurrence relation derived from the DE (5).
Within this framework, the goal of symbolic reduction is transformed into an algebraic problem: expressing a higher-degree operator as a linear combination of lower-degree operators, using the DEs derived from the fundamental IBP relations. The primary objective of this work is to develop a systematic algorithm to achieve this task.
III Symbolic Reduction Rules
The DEs derived from fundamental IBP identities can be classified based on the highest-degree operators. This classification directly informs how symbolic reduction rules are extracted:
  • I:
    The equation contains a single, unique highest degree operator, which can be directly solved to yield an immediate reduction rule.
  • II:
    The equation contains multiple highest degree operators that share the same operator index. When translated into the recurrence form (6), they effectively leave a single highest-degree term, yielding a reduction rule.
  • III:
    The equation contains multiple highest degree operators with different operator indices. In this case, one operator is chosen as the target to solve, according to the aim of reduction.
Given a reduction rule, i.e, expressing a higher operator index $\vec{o}$ by others, we can determine which points can be reduced or can not. They are given by two sets, the reduced set
${\cal S}_{\widehat{O}}=\{\vec{n}\in Z^{N}|\vec{n}\geq 0\penalty 10000\ \penalty 10000\ \&\penalty 10000\ \penalty 10000\ \vec{n}-\vec{o}\geq 0\}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$(7)
and unreduced set
${\cal U}_{\widehat{O}}=\{\vec{n}\in Z^{N}|\vec{n}\geq 0\penalty 10000\ \penalty 10000\ \&\penalty 10000\ \penalty 10000\ \exists i,(\vec{n}-\vec{o})_{i}<0\}\penalty 10000\ .\penalty 10000\ \penalty 10000\ \penalty 10000\$(8)
A crucial component of our methodology is the generation of "descendant" rules by acting on an existing DE with differential operators. The complete algorithm is structured as an iterative loop that expands the set of known reduction rules until a complete basis is found.
We now outline the core algorithm for deriving a complete set of symbolic reduction rules:
  1. (A)
    Initialization: Generate an initial set of DEs from the fundamental IBP relations. Extract immediate reduction rules and define the initial reduced set $\mathcal{S}$ and unreduced set $\mathcal{U}$. Initialize the maximum degree of differential operators to $M=2$.
  2. (B)
    Generation of New Equations: Create a new set of DEs by applying differential operators $\prod_{i}(\partial/\partial\eta_{i})^{k_{i}}$, up to the current maximum degree $M$, to the existing DEs. Apply all known reduction rules (as well as their descendant reduction rules) to new DEs to eliminate as many as possible highest-degree operators in favor of lower-degree ones, which typically refers to integrability conditions of differential operators. In this way, we get the DE for new unsolved highest operators.
  3. (C)
    Finding new reduction rules: Using the obtained DE to solve new highest operators based on the current unreduced set $\mathcal{U}$. Simplifying new reduction rules by setting terms in DE’s to zero when applying these reduction rules to the set $\mathcal{U}$.
  4. (D)
    Update and Completeness Check: Based on new found reduction rules, update the reduced set $\mathcal{S}$ and unreduced set $\mathcal{U}$. Check if the size of the updated $\mathcal{U}$ matches the number of master integrals. If so, the reduction is complete and the algorithm terminates successfully. Else, return to step (B).
  5. (E)
    Expand the maximum degree of differential operators: If steps above failed to produce any new reduction rules, the system has stagnated at the current order $M$. Increment the maximum degree $M\to M+1$, and return to step (B) to generate more complex new equations and find new rules.
Our framework offers distinct advantages over existing methods. Compared to methods based on gröbner basis [36, 37, 35, 38, 39], our approach circumvents complex non-commutative algebra. Instead, it systematically reduces the derivation of reduction rules to a more tractable linear algebra problem. Compared to heuristic algorithms [9, 10], our approach provides a general and reusable solution with a clear termination condition. This contrasts sharply with the ad-hoc nature of heuristic strategies, which lack guarantees of completeness and often fail when faced with exceptionally complex topologies.
Furthermore, our method streamlines previous generating function formalisms [41]. By employing a single generating function per sector, we avoid the expensive cost of constructing closed differential equations for a large number of master generating functions, which is achieved by Laporta’s algorithm in Ref. [41]. And by introducing a novel simplification technique on the unreduced set, we significantly accelerate the generation of recurrence rules and descendant equations, making the entire reduction process more direct and computationally tractable.
Following this general framework, we will present several examples to demonstrate the algorithm of finding complete symbolic reduction rules in the following section. Our analysis will focus exclusively on top-sector integrals, because any integral in sub-sector can be treated as a top-sector integral within a simpler sub-family, defined by the propagators with non-positive powers.
IV Example 1: the top sector of sunset diagrams
For the sunset, the propagators and ISPs are
$\mathcal{D}_{1}$ $=$ $(\ell_{1}^{2}-m_{1}^{2}),\penalty 10000\ \penalty 10000\ \penalty 10000\ \mathcal{D}_{2}=(\ell_{2}^{2}-m_{2}^{2}),$
$\mathcal{D}_{3}$ $=$ $((\ell_{1}+\ell_{2}-K)^{2}-m_{3}^{2}),$
$\mathcal{D}_{4}$ $=$ $\ell_{1}\cdot K,\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \mathcal{D}_{5}=\ell_{2}\cdot K.\penalty 10000\ \penalty 10000\ \penalty 10000\$(9)
Among $2^{3}=8$ possible generating function, only following four sectors are nonzero
$G_{111},G_{110},G_{101},G_{011}\penalty 10000\ \penalty 10000\ \penalty 10000\$(10)
where for simplicity $\mu_{4}=\mu_{5}=0$ have been implied.
There are six fundamental IBP relations for top-sector, producing six type-I equations for degree-2 operators ${\partial\over\partial\eta_{i}}{\partial\over\partial\eta_{j}}$, $i=4,5;j=1,2,3$ (see (22) to (27) for their explicit expressions). For example, the IBP relation $0=\int\prod_{i=1}^{L}\frac{\mathrm{d}^{D}\ell_{i}}{\mathrm{i}\pi^{D/2}}{\partial\over\partial\ell_{1}}\cdot\left\{\ell_{1}e^{\sum_{j=4,5}\eta_{j}s_{0}^{-1}\mathcal{D}_{j}}\prod_{i=1}^{3}{1\over\mathcal{D}_{i}-\eta_{i}s_{0}}\right\}$ produces
$0$ $=$ $\left\{2{\partial\over\partial\eta_{5}}{\partial\over\partial\eta_{3}}G_{111}\right\}-\left\{{2m_{1}^{2}\over s_{0}}{\partial\over\partial\eta_{1}}G_{111}+{f_{+-+}\over s_{0}}{\partial\over\partial\eta_{3}}G_{111}\right\}$ (11)
$+\left\{(D-3)G_{111}+\eta_{4}{\partial\over\partial\eta_{4}}G_{111}-2\eta_{1}{\partial\over\partial\eta_{1}}G_{111}\right$.
$\left.-(\eta_{3}+\eta_{1}-\eta_{2}){\partial\over\partial\eta_{3}}G_{111}\right\}$
$+\left\{-{1\over s_{0}}{\partial\over\partial\eta_{3}}G_{011}|_{\eta_{1}=0}+{1\over s_{0}}{\partial\over\partial\eta_{3}}G_{101}|_{\eta_{2}=0}\right\},\penalty 10000\ \penalty 10000\ \penalty 10000\$
where $f_{\pm\pm\pm}=\pm m_{1}^{2}\pm m_{2}^{2}\pm m_{3}^{2}-K^{2}$. The unreduced set under these six reduction rules is
${\cal U}=(0,0,0,n_{4},n_{5})\cup(n_{1},n_{2},n_{3},0,0)\penalty 10000\ .\penalty 10000\ \penalty 10000\ \penalty 10000\$(12)
Now we generate new equations from the above six equations by applying differential operators $\partial\over\partial\eta_{i}$}. Among $30$ equations, only $8$ independent new equations are kept according to the step (B), which is the consequence of integrability conditions like ${\partial\over\partial\eta_{k}}\left({\partial\over\partial\eta_{j}}{\partial\over\partial\eta_{i}}\right)-{\partial\over\partial\eta_{k}}\left({\partial\over\partial\eta_{k}}{\partial\over\partial\eta_{i}}\right)=0$. The behavior of new equations depend on the masses in a nontrivial way.
IV.1 The massless case $m_{i}=0,i=1,2,3$
Consider the set $(0,0,0,n_{4},n_{5})$, using $8$ equations we can find two degree-1 type-II equations:
$2\eta_{4}{\partial^{2}\over\partial\eta_{4}^{2}}G_{111}+2\eta_{5}{\partial\over\partial\eta_{4}}{\partial\over\partial\eta_{5}}G_{111}-(6-3D){\partial\over\partial\eta_{4}}G_{111}$ (13)
$=$ $\left\{{K^{2}(D-2)\over s_{0}}G_{111}+{K^{2}\over s_{0}}\eta_{5}{\partial\over\partial\eta_{5}}G_{111}\right$.
$\left.+{K^{2}\over s_{0}}(\eta_{4}+\eta_{5}){\partial\over\partial\eta_{4}}G_{111}\right\}+\left\{-{(K^{2})^{2}\eta_{5}\over 2s_{0}^{2}}G_{111}\right\}\penalty 10000\ ,\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$
$2\eta_{5}{\partial^{2}\over\partial\eta_{5}^{2}}G_{111}+2\eta_{4}{\partial\over\partial\eta_{4}}{\partial\over\partial\eta_{5}}G_{111}-(6-3D){\partial\over\partial\eta_{5}}G_{111}$ (14)
$=$ $\left\{{K^{2}\eta_{4}\over s_{0}}{\partial\over\partial\eta_{4}}G_{111}+{K^{2}(\eta_{5}+\eta_{4})\over s_{0}}{\partial\over\partial\eta_{5}}G_{111}\right$.
$\left.+{K^{2}(D-2)\over s_{0}}G_{111}\right\}+\left\{-{(K^{2})^{2}\eta_{4}\over 2s_{0}^{2}}G_{111}\right\}\penalty 10000\ .\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$
Using above two reduction rules we can reduce any point $(0,0,0,n_{4},n_{5})$ down to $(0,0,0,0,0)$.
Now consider the reduction of the set $(n_{1},n_{2},n_{3},0,0)$, we can form two degree-1 type-III equations:
$\frac{D-4}{2}{\partial\over\partial\eta_{i}}G_{111}-\eta_{i}{\partial^{2}\over\partial\eta_{i}^{2}}G_{111}$ (15)
$=$ $\frac{D-4}{2}{\partial\over\partial\eta_{3}}G_{111}-\eta_{3}{\partial^{2}\over\partial\eta_{3}^{2}}G_{111}+\mathcal{R},\ i=1,2\penalty 10000\ \penalty 10000\ \penalty 10000\$
where $\mathcal{R}$ represents lower degree terms. Under the reduction rules produced, the only unreduced set is $(0,0,n_{3},0,0)$. To finish the reduction, we need to generate another equation, which can be achieved by $\partial\over\partial\eta_{3}$} on (13) (or (14)). After simplification we will get
${(D-4)K^{2}\over s_{0}}{\partial\over\partial\eta_{3}}G_{111}-{2K^{2}\over s_{0}}\eta_{3}{\partial^{2}\over\partial\eta_{3}^{2}}G_{111}=-2\eta_{3}^{2}{\partial^{2}\over\partial\eta_{3}^{2}}G_{111}$
${(5D-16)\eta_{3}}{\partial\over\partial\eta_{3}}G_{111}-{(D-3)(3D-8)}G_{111}.\penalty 10000\ \penalty 10000\ \penalty 10000\$(16)
All nonzero $n_{3}$ can be reduced to $n_{3}=0$ by the reduction rule. Eventually, the only point can not be reduced is $(0,0,0,0,0)$, which is the master integral for the massless case.
IV.2 The massive case
The reduction of the set $(0,0,0,n_{4},n_{5})$ is similar to the massless case, i.e., we can construct equations like (13) and (14). However, to reduce the set $(n_{1},n_{2},n_{3},0,0)$, there are only $5$ independent degree-2 equations with six degree operators ${\partial^{2}\over\partial\eta_{i}^{2}},i=1,2,3$ and ${\partial\over\partial\eta_{i}}{\partial\over\partial\eta_{j}},1\leq i<j\leq 3$. Again, after acting $\partial\over\partial\eta_{1}$} on (13), we can form the sixth equation to solve six degree two operators. The final result is that there are four points can not be reduced, i.e., (0,0,0,0,0),(1,0,0,0,0),(0,1,0,0,0),(0,0,1,0,0)(0,0,0,0,0),(1,0,0,0,0),(0,1,0,0,0),(0,0,1,0,0), which give the master integrals for massive sunset.
V Example 2: top sector of two-loop massless double-box
For this one, four external momenta $k_{1}$, $k_{2}$, $k_{3}$, $k_{4}$ satisfy on-shell conditions $k_{i}^{2}=0$ and momentum conservation $\sum_{i=1}^{4}k_{i}=0$, which leaves two independent scales $s=2k_{1}\cdot k_{2}$ and $t=2k_{2}\cdot k_{3}$. We choose a complete set of Lorentz scalars as
$\begin{split}&\mathcal{D}_{1}={\ell_{1}}^{2},\ \mathcal{D}_{2}=(\ell_{1}+k_{1})^{2},\ \mathcal{D}_{3}={(\ell_{1}+k_{1}+k_{2})}^{2},\\ &\mathcal{D}_{4}={\ell_{2}}^{2},\ \mathcal{D}_{5}={(\ell_{2}-\ell_{1})}^{2},\ \mathcal{D}_{6}={(\ell_{2}+k_{1}+k_{2})}^{2},\\ &\mathcal{D}_{7}={(\ell_{2}-k_{4})}^{2},\ \mathcal{D}_{8}=\ell_{1}\cdot k_{3},\ \mathcal{D}_{9}=\ell_{2}\cdot k_{1},\penalty 10000\ \penalty 10000\ \penalty 10000\ \end{split}$ (17)
where the last two are ISPs.
There are $10$ fundamental IBP relations. From them, we can solve following $4$ degree-1 operators ${\partial\over\partial\eta_{i}},i=1,3,4,6$ and $6$ degree-2 operators ${\partial\over\partial\eta_{i}}{\partial\over\partial\eta_{j}},i=2,5,7;j=8,9$. The unreduced sets under these $10$ reduction rules are
${\cal U}_{1}$ $=$ $(0,n_{2},0,0,n_{5},0,n_{7},0,0),$
${\cal U}_{2}$ $=$ $(0,0,0,0,0,0,0,n_{8},n_{9})\penalty 10000\ \penalty 10000\ \penalty 10000\ ,$ (18)
which are similar with the cases of top-sector of sunset. By the step (B), We can generate $8$ independent equations by applying differential operators to $6$ degree-2 equations.
Three of them, i.e., ${\partial\over\partial\eta_{9}}\left({\partial\over\partial\eta_{8}}{\partial\over\partial\eta_{i}}\right)-{\partial\over\partial\eta_{8}}\left({\partial\over\partial\eta_{9}}{\partial\over\partial\eta_{i}}\right)=0,i=2,5,7$, can be applied to produce recurrence relation for $F_{\vec{n}+\vec{e}_{8}+\vec{e}_{9}},F_{\vec{n}+\vec{e}_{8}},F_{\vec{n}+\vec{e}_{9}}$, which reduce the set ${\cal U}_{2}$ to point $n_{8}=n_{9}=0$. Considering other $5$ equations for reducing the set ${\cal U}_{1}$, $3$ of them are degree-2 type-I equations for operators ${\partial\over\partial\eta_{2}}{\partial\over\partial\eta_{5}},{\partial\over\partial\eta_{5}}{\partial\over\partial\eta_{7}},{\partial\over\partial\eta_{2}}{\partial\over\partial\eta_{7}}$. Another two give degree-1 type-III equations
${{D-6}\over 2}{\partial\over\partial\eta_{i}}G-\eta_{i}{\partial^{2}\over\partial\eta_{i}^{2}}G$ (19)
$=$ ${{D-6}\over 2}{\partial\over\partial\eta_{5}}G-\eta_{5}{\partial^{2}\over\partial\eta_{5}^{2}}G+\mathcal{R},\ i=2,7\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$
Iteratively using reducing rules from these equation, the unreduced set is just $(0,0,0,0,n_{5},0,0,0,0)$. By acting $\partial\over\partial\eta_{5}$} on (19) with $i=2$, and using the reduction rule of ${\partial\over\partial\eta_{2}}{\partial\over\partial\eta_{5}}$ just found, we can get a recurrence relation for $F_{\vec{n}+2\vec{e}_{5}}$, which reduce any points of $n_{5}\geq 2$. Now the only points can not be reduced is $n_{5}=0,1$, which are two master integrals for top-sector of massless double-box.
VI Example 3: top sector of two-loop massless non-planer
The kinematics are the same as planer double-box. We choose a complete set of Lorentz scalars as
$\begin{split}&\mathcal{D}_{1}={\ell_{1}}^{2},\penalty 10000\ \penalty 10000\ \mathcal{D}_{2}=(\ell_{1}-k_{1})^{2},\penalty 10000\ \penalty 10000\ \penalty 10000\ \mathcal{D}_{3}={(\ell_{1}-k_{1}-k_{2})}^{2},\\ &\mathcal{D}_{4}={\ell_{2}}^{2},\penalty 10000\ \mathcal{D}_{5}={(\ell_{2}+k_{4})}^{2},\penalty 10000\ \mathcal{D}_{6}={(\ell_{2}-\ell_{1}+k_{1}+k_{2}+k_{4})}^{2},\\ &\mathcal{D}_{7}={(\ell_{2}-\ell_{1})}^{2},\penalty 10000\ \penalty 10000\ \penalty 10000\ \mathcal{D}_{8}=\ell_{1}\cdot k_{4},\penalty 10000\ \penalty 10000\ \penalty 10000\ \mathcal{D}_{9}=\ell_{2}\cdot k_{1},\penalty 10000\ \penalty 10000\ \penalty 10000\ \end{split}$ (20)
where the last two are ISPs.
There are $10$ fundamental IBP relations. Using them we can solve two degree-1 operators $\partial\over\partial\eta_{1}$}, $\partial\over\partial\eta_{3}$} and $6$ degree-2 operators ${\partial\over\partial\eta_{i}}{\partial\over\partial\eta_{8}},i=2,4,5,6,7$ and ${\partial\over\partial\eta_{2}}{\partial\over\partial\eta_{9}}$. There are still four unsolved degree-2 operators ${\partial\over\partial\eta_{i}}{\partial\over\partial\eta_{9}},i=4,5,6,7$, which appear in two equations in group form, i.e., $\left({\partial\over\partial\eta_{4}}+{\partial\over\partial\eta_{5}}\right){\partial\over\partial\eta_{9}}$ and $\left({\partial\over\partial\eta_{6}}+{\partial\over\partial\eta_{7}}\right){\partial\over\partial\eta_{9}}$. Adding two new integrability conditions, i.e., ${\partial\over\partial\eta_{2}}\left({\partial\over\partial\eta_{4}}+{\partial\over\partial\eta_{5}}\right){\partial\over\partial\eta_{9}}-\left({\partial\over\partial\eta_{4}}+{\partial\over\partial\eta_{5}}\right){\partial\over\partial\eta_{2}}{\partial\over\partial\eta_{9}}=0$ and ${\partial\over\partial\eta_{2}}\left({\partial\over\partial\eta_{6}}+{\partial\over\partial\eta_{7}}\right){\partial\over\partial\eta_{9}}-\left({\partial\over\partial\eta_{6}}+{\partial\over\partial\eta_{7}}\right){\partial\over\partial\eta_{2}}{\partial\over\partial\eta_{9}}=0$, we can solve these four operators. Using these reduction rules, now the unreduced sets becomes
${\cal U}_{1}$ $=$ $(0,n_{2},0,n_{4},n_{5},n_{6},n_{7},0,0),$
${\cal U}_{2}$ $=$ $(0,0,0,0,0,0,0,n_{8},n_{9})$ (21)
which becomes the familiar pattern, like the one meet in the top sector of sunset.
Now we do similar procedure. Using the new equations coming from integrability condition ${\partial\over\partial\eta_{9}}\left({\partial\over\partial\eta_{i}}{\partial\over\partial\eta_{8}}\right)-{\partial\over\partial\eta_{8}}\left({\partial\over\partial\eta_{i}}{\partial\over\partial\eta_{9}}\right)=0,i=2,4,5,6,7$, we can reduce the set ${\cal U}_{2}$ to point $n_{8}=n_{9}=0$. Using the new equations coming from integrability condition ${\partial\over\partial\eta_{j}}\left({\partial\over\partial\eta_{i}}{\partial\over\partial\eta_{k}}\right)-{\partial\over\partial\eta_{i}}\left({\partial\over\partial\eta_{j}}{\partial\over\partial\eta_{k}}\right)=0,i<j,i,j=2,4,5,6,7;k=8,9$, we can solve degree two operators ${\partial\over\partial\eta_{j}}{\partial\over\partial\eta_{i}}$ and four degree one operator ${\partial\over\partial\eta_{i}},i=2,5,6,7$. Finally by generating a new degree two equation for $\partial^{2}\over\partial\eta_{4}^{2}$}, we are able to reduce the set ${\cal U}_{1}$ to two points $n_{2}=n_{5}=n_{6}=n_{7}=0$, $n_{4}=0,1$, i.e., two master integrals. For more details, see the supplemental material.
VII Conclusion and Outlook
In this work, we introduced a novel systematic approach to derive complete symbolic reduction rules for multi-loop Feynman integrals with generating functions. By the DEs for these generating functions, we obtain recurrence relations that enable the efficient reduction without the exponential proliferation of IBP identities in traditional methods. Through explicity examples, including the sunset diagrams (both massive and massless), two-loop double-box, and two-loop non-planar integrals, we demonstrate the algorithm’s effectiveness in complex cases with high-degree numerators and high-power propagators. This method addresses key bottlenecks in perturbative QFT computations.
With the proof of concept in the note, the most important thing is to automate the aforementioned algorithm. In a longer companion paper, more details of the algorithm will be presented. The combination of the idea in the paper and other reduction methods is also worth pursuing.
Acknowledgements.
We thank Mingxing Luo for many helpfull discussions. Bo Feng is supported by the National Natural Science Foundation of China (NSFC) through Grants No. 12535003, No.11935013, No.11947301, No.12047502. Xiang Li and Yan-Qing Ma are supported by NSFC through Grant No. 12325503. Yuanche Liu is supported by NSFC through Grant No. 124B1014, and Yang Zhang is supported by NSFC through Grant No. 12575078 and 12247103.
Supplemental Material
In the supplemental material, we provide some details of the examples in this paper.
Appendix A The topsector of sunset
Using six DEs coming from six fundamental IBP relations, we can solve six degree two operators as
$2{\partial\over\partial\eta_{5}}{\partial\over\partial\eta_{3}}G_{111}=-\left\{{2m_{1}^{2}\over s_{0}}{\partial\over\partial\eta_{1}}G_{111}+{f_{+-+}\over s_{0}}{\partial\over\partial\eta_{3}}G_{111}\right\}$
$+\left\{(D-3)G_{111}+\eta_{4}{\partial\over\partial\eta_{4}}G_{111}-2\eta_{1}{\partial\over\partial\eta_{1}}G_{111}\right$.
$\left.-(\eta_{3}+\eta_{1}-\eta_{2}){\partial\over\partial\eta_{3}}G_{111}\right\}\penalty 10000\ \penalty 10000\ \penalty 10000\$(22)
$2{\partial\over\partial\eta_{5}}{\partial\over\partial\eta_{1}}G_{111}=\left\{-{f_{+-+}\over s_{0}}{\partial\over\partial\eta_{1}}G_{111}-{2m_{3}^{2}\over s_{0}}{\partial\over\partial\eta_{3}}G_{111}\right\}$
$+\left\{(D-3)G_{111}+\eta_{4}{\partial\over\partial\eta_{4}}G_{111}+\eta_{4}{\partial\over\partial\eta_{5}}G_{111}\right$.
$\left.+(-\eta_{3}-\eta_{1}+\eta_{2}){\partial\over\partial\eta_{1}}G_{111}-2\eta_{3}{\partial\over\partial\eta_{3}}G_{111}\right\}$
$+\left\{-\eta_{4}K^{2}s_{0}^{-1}G_{111}\right\}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$(23)
$2{\partial\over\partial\eta_{4}}{\partial\over\partial\eta_{1}}G_{111}=\left\{{2m_{3}^{2}\over s_{0}}{\partial\over\partial\eta_{3}}G_{111}+{2m_{1}^{2}\over s_{0}}{\partial\over\partial\eta_{1}}G_{111}\right$.
$\left.+{2m_{2}^{2}\over s_{0}}{\partial\over\partial\eta_{2}}G_{111}\right\}+\left\{-2(D-3)G_{111}-\eta_{4}{\partial\over\partial\eta_{4}}G_{111}\right$.
$\left.-\eta_{5}{\partial\over\partial\eta_{5}}G_{111}+2\eta_{1}{\partial\over\partial\eta_{1}}G_{111}+2\eta_{2}{\partial\over\partial\eta_{2}}G_{111}\right$.
$\left.+2\eta_{3}{\partial\over\partial\eta_{3}}G_{111}\right\}+\left\{\eta_{4}K^{2}s_{0}^{-1}G_{111}\right\}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$(24)
$2{\partial\over\partial\eta_{4}}{\partial\over\partial\eta_{2}}G_{111}=\left\{-{f_{-++}\over s_{0}}{\partial\over\partial\eta_{2}}G_{111}-{2m_{3}^{2}\over s_{0}}{\partial\over\partial\eta_{3}}G_{111}\right\}$
$+\left\{(D-3)G_{111}+\eta_{5}{\partial\over\partial\eta_{4}}G_{111}+\eta_{5}{\partial\over\partial\eta_{5}}G_{111}\right$.
$\left.+(-\eta_{3}+\eta_{1}-\eta_{2}){\partial\over\partial\eta_{2}}G_{111}-2\eta_{3}{\partial\over\partial\eta_{3}}G_{111}\right\}$
$+\left\{-\eta_{5}K^{2}s_{0}^{-1}G_{111}\right\}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$(25)
$2{\partial\over\partial\eta_{4}}{\partial\over\partial\eta_{3}}G_{111}=-\left\{{2m_{2}^{2}\over s_{0}}{\partial\over\partial\eta_{2}}G_{111}+{f_{-++}\over s_{0}}{\partial\over\partial\eta_{3}}G_{111}\right\}$
$+\left\{(D-3)G_{111}+\eta_{5}{\partial\over\partial\eta_{5}}G_{111}-2\eta_{2}{\partial\over\partial\eta_{2}}G_{111}\right$.
$\left.-(\eta_{3}-\eta_{1}+\eta_{2}){\partial\over\partial\eta_{3}}G_{111}\right\}\penalty 10000\ \penalty 10000\ \penalty 10000\$(26)
$2{\partial\over\partial\eta_{5}}{\partial\over\partial\eta_{2}}G_{111}=\left\{{2m_{3}^{2}\over s_{0}}{\partial\over\partial\eta_{3}}G_{111}+{2m_{1}^{2}\over s_{0}}{\partial\over\partial\eta_{1}}G_{111}\right$.
$\left.+{2m_{2}^{2}\over s_{0}}{\partial\over\partial\eta_{2}}G_{111}\right\}\ +\left\{-2(D-3)G_{111}-\eta_{4}{\partial\over\partial\eta_{4}}G_{111}\right$.
$\left.-\eta_{5}{\partial\over\partial\eta_{5}}G_{111}+2\eta_{1}{\partial\over\partial\eta_{1}}G_{111}+2\eta_{2}{\partial\over\partial\eta_{2}}G_{111}\right$.
$\left.+2\eta_{3}{\partial\over\partial\eta_{3}}G_{111}\right\}+\left\{\eta_{5}K^{2}s_{0}^{-1}G_{111}\right\}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$(27)
where since we focus on the top sector only, contributions from subsectors have been neglected. Also, we have written into the form with degree separated. They give six good reduction rules. Using them, we can reduce any point except the following two sets ${\cal U}_{1}=(n_{1},n_{2},n_{3},0,0)$ and ${\cal U}_{2}=(0,0,0,n_{4},n_{5})$. To reduce the two remaining sets, we need to generate new equations. Acting on $\partial\over\partial\eta_{i}$} on above six equations, we will get $30$ degree $3$ equations. However, by using descendant reduction rules of above six reduction rules, one can check that we get $9$ degree $2$ equations, where one of them is
$0$ $=$ $\left\{-\eta_{4}{\partial^{2}\over\partial\eta_{4}^{2}}G_{111}-\eta_{5}{\partial^{2}\over\partial\eta_{5}^{2}}G_{111}-2\eta_{4}{\partial\over\partial\eta_{4}}{\partial\over\partial\eta_{5}}G_{111}\right$. (28)
$\left.+{2m_{1}^{2}(m_{2}^{2}-m_{3}^{2})\over s_{0}^{2}}{\partial\over\partial\eta_{1}}G_{111}+{m_{2}^{2}f_{++-}\over s_{0}^{2}}{\partial\over\partial\eta_{2}}G_{111}\right$.
$\left.-{m_{3}^{2}f_{+-+}\over s_{0}^{2}}{\partial\over\partial\eta_{3}}G_{111}-(D-2){\partial\over\partial\eta_{4}}G_{111}\right$.
$\left.+(4-2D){\partial\over\partial\eta_{5}}G_{111}\right\}$
$+\left\{+{2(m_{2}^{2}-m_{3}^{2})\eta_{1}+2m_{1}^{2}(\eta_{2}-\eta_{3})\over s_{0}}{\partial\over\partial\eta_{1}}G_{111}\right$.
$\left.-{m_{2}^{2}(\eta_{1}+\eta_{2}-\eta_{3})+f_{++-}\eta_{2}\over s_{0}}{\partial\over\partial\eta_{2}}G_{111}\right$.
$+{-f_{+-+}\eta_{3}-m_{3}^{2}(\eta_{1}-\eta_{2}+\eta_{3})\over s_{0}}{\partial\over\partial\eta_{3}}G_{111}$
$+{2K^{2}-f_{-+-}\over 2s_{0}}\eta_{4}{\partial\over\partial\eta_{4}}G_{111}$
$\left.+{2(K^{2}+m_{1}^{2})\eta_{4}-f_{++-}\eta_{5}\over 2s_{0}}{\partial\over\partial\eta_{5}}G_{111}\right$.
$\left.+{(D-2)K^{2}-(D-3)(m_{2}^{2}-m_{3}^{2})\over s_{0}}G_{111}\right\}$
$+\left\{+2(\eta_{2}-\eta_{3})\eta_{1}{\partial\over\partial\eta_{1}}G_{111}+(\eta_{1}+\eta_{2}-\eta_{3})\eta_{2}{\partial\over\partial\eta_{2}}G_{111}\right$.
$\left.-(\eta_{1}-\eta_{2}+\eta_{3})\eta_{3}{\partial\over\partial\eta_{3}}G_{111}+{(\eta_{1}-\eta_{2}+\eta_{3})\over 2}\eta_{4}{\partial\over\partial\eta_{4}}G_{111}\right$.
$+(\eta_{1}\eta_{4}-{1\over 2}\eta_{5}(\eta_{1}+\eta_{2}-\eta_{3})){\partial\over\partial\eta_{5}}G_{111}$
$\left.+({K^{2}f_{--+}\over 2s_{0}^{2}}\eta_{4}+{K^{2}m_{2}^{2}\over s_{0}^{2}}\eta_{5}-(D-3)(\eta_{2}-\eta_{3}))G_{111}\right\}$
$+\left\{+({K^{2}\over s_{0}}\eta_{2}\eta_{5}-{K^{2}\over 2s_{0}}\eta_{4}(\eta_{1}+\eta_{2}-\eta_{3}))G_{111}\right\}\penalty 10000\ \penalty 10000\ \penalty 10000\$
These equations are, in fact, the consequence of integrability conditions ${\partial\over d\eta_{i}},{\partial\over\partial\eta_{j}}]=0$. Above $9$ new equations can be divided into three groups. The first group is given by ${\partial\over\partial\eta_{4}}\left({\partial\over\partial\eta_{5}}{\partial\over\partial\eta_{i}}\right)-{\partial\over\partial\eta_{5}}\left({\partial\over\partial\eta_{4}}{\partial\over\partial\eta_{i}}\right)=0$ with $i=1,2,3$. We will denote the corresponding equations as (I-i) ((28) is (I-1)). The second group is given by ${\partial\over\partial\eta_{j}}\left({\partial\over\partial\eta_{4}}{\partial\over\partial\eta_{i}}\right)-{\partial\over\partial\eta_{i}}\left({\partial\over\partial\eta_{4}}{\partial\over\partial\eta_{j}}\right)=0$ with $1\leq i<j\leq 3$, which will be denoted as (II-ij). The third group is given by ${\partial\over\partial\eta_{j}}\left({\partial\over\partial\eta_{5}}{\partial\over\partial\eta_{i}}\right)-{\partial\over\partial\eta_{i}}\left({\partial\over\partial\eta_{5}}{\partial\over\partial\eta_{j}}\right)=0$ with $1\leq i<j\leq 3$, which will be denoted as (III-ij). These $9$ equations are not independent since  ${\text{ (II-12)}}+{\text{ (III-12)}}-{\text{ (II-23)}}-{\text{ (III-13)}}=0$. Also the behavior of new equations depend on the masses in a nontrivial way.
The result (28) is presented according to the degree of operators. For later discussions, only the degree one and two operators are important, so we will neglect the other parts. Similarly, the relevant part of (I-2) is
$0$ $=$ $\left\{-\eta_{4}{\partial^{2}\over\partial\eta_{4}^{2}}G_{111}-\eta_{5}{\partial^{2}\over\partial\eta_{5}^{2}}G_{111}-2\eta_{5}{\partial\over\partial\eta_{4}}{\partial\over\partial\eta_{5}}G_{111}\right$. (29)
$\left.+{2m_{2}^{2}(m_{1}^{2}-m_{3}^{2})\over s_{0}^{2}}{\partial\over\partial\eta_{2}}G_{111}+{m_{1}^{2}f_{++-}\over s_{0}^{2}}{\partial\over\partial\eta_{1}}G_{111}\right$.
$\left.-{m_{3}^{2}f_{-++}\over s_{0}^{2}}{\partial\over\partial\eta_{3}}G_{111}+(4-2D){\partial\over\partial\eta_{4}}G_{111}\right$.
$\left.+(2-D){\partial\over\partial\eta_{5}}G_{111}\right\}+.$$penalty 10000\ \penalty 10000\$
and the relevant part of (I-3) is
$0$ $=$ $\left\{-\eta_{4}{\partial^{2}\over\partial\eta_{4}^{2}}G_{111}+\eta_{5}{\partial^{2}\over\partial\eta_{5}^{2}}G_{111}+{m_{1}^{2}f_{+-+}\over s_{0}^{2}}{\partial\over\partial\eta_{1}}G_{111}\right$. (30)
$-{m_{2}^{2}f_{-++}\over s_{0}^{2}}{\partial\over\partial\eta_{2}}G_{111}+{2m_{3}^{2}(m_{1}^{2}-m_{2}^{2})\over s_{0}^{2}}{\partial\over\partial\eta_{3}}G_{111}$
$\left.-(D-2){\partial\over\partial\eta_{4}}G_{111}+(D-2){\partial\over\partial\eta_{5}}G_{111}\right\}+$$penalty 10000\ \penalty 10000\$
The special point of the above three equations is that the highest operators have degree one. Groups II and III are degree two equations. The (II-12) is
${2m_{2}^{2}\over s_{0}}{\partial^{2}\over\partial\eta_{2}^{2}}G_{111}+{f_{+++}\over s_{0}}{\partial\over\partial\eta_{2}}{\partial\over\partial\eta_{1}}G_{111}$ (31)
$+{2m_{3}^{2}\over s_{0}}{\partial\over\partial\eta_{3}}{\partial\over\partial\eta_{1}}G_{111}+{2m_{3}^{2}\over s_{0}}{\partial\over\partial\eta_{3}}{\partial\over\partial\eta_{2}}G_{111}$
$=$ $\left\{-2\eta_{2}{\partial^{2}\over\partial\eta_{2}^{2}}G_{111}-(\eta_{1}+\eta_{2}+\eta_{3}){\partial\over\partial\eta_{2}}{\partial\over\partial\eta_{1}}G_{111}\right$.
$\left.-2\eta_{3}{\partial\over\partial\eta_{3}}{\partial\over\partial\eta_{1}}G_{111}-2\eta_{3}{\partial\over\partial\eta_{3}}{\partial\over\partial\eta_{2}}G_{111}\right$.
$\left.+(D-3){\partial\over\partial\eta_{1}}G_{111}+(2D-7){\partial\over\partial\eta_{2}}G_{111}\right\}+$$penalty 10000\ \penalty 10000\$
The (II-13) is
${2m_{2}^{2}\over s_{0}}{\partial\over\partial\eta_{3}}{\partial\over\partial\eta_{2}}G_{111}+{2m_{2}^{2}\over s_{0}}{\partial\over\partial\eta_{2}}{\partial\over\partial\eta_{1}}G_{111}$ (32)
$+{f_{+++}\over s_{0}}{\partial\over\partial\eta_{3}}{\partial\over\partial\eta_{1}}G_{111}+{2m_{3}^{2}\over s_{0}}{\partial^{2}\over\partial\eta_{3}^{2}}G_{111}$
$=$ $\left\{(D-3){\partial\over\partial\eta_{1}}G_{111}+(2D-7){\partial\over\partial\eta_{3}}G_{111}-2\eta_{3}{\partial^{2}\over\partial\eta_{3}^{2}}G_{111}\right$.
$\left.-2\eta_{2}{\partial\over\partial\eta_{3}}{\partial\over\partial\eta_{2}}G_{111}-2\eta_{2}{\partial\over\partial\eta_{2}}{\partial\over\partial\eta_{1}}G_{111}\right$.
$\left.-(\eta_{1}+\eta_{2}+\eta_{3}){\partial\over\partial\eta_{3}}{\partial\over\partial\eta_{1}}G_{111}\right\}+$$penalty 10000\ \penalty 10000\$
The (II-23) is
${2m_{2}^{2}\over s_{0}}{\partial^{2}\over\partial\eta_{2}^{2}}G_{111}-{2m_{3}^{2}\over s_{0}}{\partial^{2}\over\partial\eta_{3}^{2}}G_{111}$ (33)
$=$ $\left\{-(D-4){\partial\over\partial\eta_{3}}G_{111}+(D-4){\partial\over\partial\eta_{2}}G_{111}+2\eta_{3}{\partial^{2}\over\partial\eta_{3}^{2}}G_{111}\right$.
$\left.-2\eta_{2}{\partial^{2}\over\partial\eta_{2}^{2}}G_{111}\right\}+$$penalty 10000\ \penalty 10000\$
The (III-12) is
${-2m_{1}^{2}\over s_{0}}{\partial^{2}\over\partial\eta_{1}^{2}}G_{111}-{f_{+++}\over s_{0}}{\partial\over\partial\eta_{2}}{\partial\over\partial\eta_{1}}G_{111}$ (34)
$-{2m_{3}^{2}\over s_{0}}{\partial\over\partial\eta_{3}}{\partial\over\partial\eta_{1}}G_{111}-{2m_{3}^{2}\over s_{0}}{\partial\over\partial\eta_{3}}{\partial\over\partial\eta_{2}}G_{111}$
$=$ $\left\{2\eta_{1}{\partial^{2}\over\partial\eta_{1}^{2}}G_{111}+(\eta_{1}+\eta_{2}+\eta_{3}){\partial\over\partial\eta_{2}}{\partial\over\partial\eta_{1}}G_{111}\right$.
$\left.+2\eta_{3}{\partial\over\partial\eta_{3}}{\partial\over\partial\eta_{1}}G_{111}+2\eta_{3}{\partial\over\partial\eta_{3}}{\partial\over\partial\eta_{2}}G_{111}\right$.
$\left.+(7-2D){\partial\over\partial\eta_{1}}G_{111}-(D-3){\partial\over\partial\eta_{2}}G_{111}\right\}+$$penalty 10000\ \penalty 10000\$
The (III-13) is
$-{2m_{1}^{2}\over s_{0}}{\partial^{2}\over\partial\eta_{1}^{2}}G_{111}+{2m_{3}^{2}\over s_{0}}{\partial^{2}\over\partial\eta_{3}^{2}}G_{111}$ (35)
$=$ $\left\{(D-4){\partial\over\partial\eta_{3}}G_{111}-(D-4){\partial\over\partial\eta_{1}}G_{111}-2\eta_{3}{\partial^{2}\over\partial\eta_{3}^{2}}G_{111}\right$.
$\left.+2\eta_{1}{\partial^{2}\over\partial\eta_{1}^{2}}G_{111}\right\}+$$penalty 10000\ \penalty 10000\$
The (III-23) is
$-{f_{+++}\over s_{0}}{\partial\over\partial\eta_{3}}{\partial\over\partial\eta_{2}}G_{111}-{2m_{1}^{2}\over s_{0}}{\partial\over\partial\eta_{2}}{\partial\over\partial\eta_{1}}G_{111}$ (36)
$-{2m_{1}^{2}\over s_{0}}{\partial\over\partial\eta_{3}}{\partial\over\partial\eta_{1}}G_{111}-{2m_{3}^{2}\over s_{0}}{\partial^{2}\over\partial\eta_{3}^{2}}G_{111}$
$=$ $\left\{(7-2D){\partial\over\partial\eta_{3}}G_{111}-(D-3){\partial\over\partial\eta_{2}}G_{111}+2\eta_{3}{\partial^{2}\over\partial\eta_{3}^{2}}G_{111}\right$.
$\left.+(\eta_{1}+\eta_{2}+\eta_{3}){\partial\over\partial\eta_{3}}{\partial\over\partial\eta_{2}}G_{111}+2\eta_{1}{\partial\over\partial\eta_{2}}{\partial\over\partial\eta_{1}}G_{111}\right$.
$\left.+2\eta_{1}{\partial\over\partial\eta_{3}}{\partial\over\partial\eta_{1}}G_{111}\right\}+$$penalty 10000\ \penalty 10000\$
A crucial point for the group II and III is that when we consider the massless limit, some degree two relations will be reduced to the degree one relations, such as (33) and (35). This difference will lead to different master integrals, which have been discussed in the main text.
To finish the reduction of the set ${\cal U}_{1}$ for massive case, we need to generate one more equation, which comes from ${\partial\over\partial\eta_{1}}({\text{(I-2)+(I-3)}})$ and is given by
${2m_{1}^{2}(K^{2}-m_{1}^{2})\over s_{0}^{2}}{\partial^{2}\over\partial\eta_{1}^{2}}G_{111}$ (37)
$+{m_{2}^{2}(f_{-++}-2(m_{1}^{2}-m_{3}^{2}))\over s_{0}^{2}}{\partial\over\partial\eta_{2}}{\partial\over\partial\eta_{1}}G_{111}$
$+{m_{3}^{2}(f_{-++}-2(m_{1}^{2}-m_{2}^{2}))\over s_{0}^{2}}{\partial\over\partial\eta_{3}}{\partial\over\partial\eta_{1}}G_{111}$
$=$ $\left\{{-3(D-3)m_{3}^{2}\over s_{0}}{\partial\over\partial\eta_{3}}G_{111}+{-3(D-3)m_{2}^{2}\over s_{0}}{\partial\over\partial\eta_{2}}G_{111}\right$.
$-{2(K^{2}-2m_{1}^{2})\over s_{0}}\eta_{1}{\partial^{2}\over\partial\eta_{1}^{2}}G_{111}$
$+\left({(D-3)(m_{2}^{2}+m_{3}^{2})+(D-4)K^{2}\over s_{0}}\right$.
$\left.+{(16-5D)m_{1}^{2}\over s_{0}}\right){\partial\over\partial\eta_{1}}G_{111}$
$+\left({(-f_{-+-}+2(m_{1}^{2}-m_{2}^{2}))\eta_{3}\over s_{0}}\right$.
$\left.+{3m_{3}^{2}(\eta_{1}-\eta_{2}-\eta_{3})\over s_{0}}\right){\partial\over\partial\eta_{3}}{\partial\over\partial\eta_{1}}G_{111}$
$+\left({(-f_{--+}+2(m_{1}^{2}-m_{3}^{2}))\eta_{2}\over s_{0}}\right$.
$\left.\left.+{3m_{2}^{2}(\eta_{1}-\eta_{2}-\eta_{3})\over s_{0}}\right){\partial\over\partial\eta_{2}}{\partial\over\partial\eta_{1}}G_{111}\right\}+.$...
where again only nonzero contribution terms for reducing the set $(n_{1},n_{2},n_{3},0,0)$ have been kept.
Appendix B The top sector of massless double box
From $10$ fundamental IBP relations, we can solve the following $4$ degree one operators
${s\over s_{0}}{\partial\over\partial\eta_{3}}G=-\left\{(D-5)G-2\eta_{1}{\partial\over\partial\eta_{1}}G-(\eta_{1}+\eta_{2}){\partial\over\partial\eta_{2}}G\right$.
$\left.-(\eta_{1}+\eta_{3}){\partial\over\partial\eta_{3}}G+(\eta_{4}-\eta_{1}-\eta_{5}){\partial\over\partial\eta_{5}}G+\eta_{8}{\partial\over\partial\eta_{8}}G\right\}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$(38)
${s\over s_{0}}{\partial\over\partial\eta_{6}}G=-\left\{(D-5)G+(\eta_{1}-\eta_{4}-\eta_{5}){\partial\over\partial\eta_{5}}G+\eta_{9}{\partial\over\partial\eta_{9}}G\right$.
$\left.-2\eta_{4}{\partial\over\partial\eta_{4}}G-(\eta_{4}+\eta_{6}){\partial\over\partial\eta_{6}}G-(\eta_{4}+\eta_{7}){\partial\over\partial\eta_{7}}G\right\}\penalty 10000\ \penalty 10000\$(39)
${s\over s_{0}}{\partial\over\partial\eta_{1}}G=-\left\{(D-5)G-(\eta_{1}+\eta_{3}){\partial\over\partial\eta_{1}}G-(\eta_{2}+\eta_{3}){\partial\over\partial\eta_{2}}G\right$.
$\left.-2\eta_{3}{\partial\over\partial\eta_{3}}G+(\eta_{6}-\eta_{3}-\eta_{5}){\partial\over\partial\eta_{5}}G+\eta_{8}{\partial\over\partial\eta_{8}}G\right\}$
$+\left\{{s\over 2s_{0}}\eta_{8}G\right\}\penalty 10000\$(40)
${s\over s_{0}}{\partial\over\partial\eta_{4}}G=\left\{-(D-5)G+(\eta_{6}+\eta_{4}){\partial\over\partial\eta_{4}}G\right$.
$\left.+(\eta_{6}-\eta_{3}+\eta_{5}){\partial\over\partial\eta_{5}}G+(\eta_{6}+\eta_{7}){\partial\over\partial\eta_{7}}G\right$.
$\left.+2\eta_{6}{\partial\over\partial\eta_{6}}G-\eta_{9}{\partial\over\partial\eta_{9}}G\right\}+\left\{{-s\over 2s_{0}}\eta_{9}G\right\}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$(41)
and $6$ degree two operators
$2{\partial\over\partial\eta_{9}}{\partial\over\partial\eta_{5}}G=-\left\{(D-5)G-(\eta_{1}+\eta_{2}){\partial\over\partial\eta_{1}}G\right$.
$\left.-2\eta_{2}{\partial\over\partial\eta_{2}}G-(\eta_{2}+\eta_{3}){\partial\over\partial\eta_{3}}G+(\eta_{4}-\eta_{2}-\eta_{5}){\partial\over\partial\eta_{5}}G\right$.
$\left.+\eta_{8}{\partial\over\partial\eta_{8}}G\right\}+\left\{{(s+t)\over 2s_{0}}\eta_{8}G\right\}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$(42)
$2{\partial\over\partial\eta_{8}}{\partial\over\partial\eta_{5}}G=\left\{{s\over s_{0}}{\partial\over\partial\eta_{5}}G\right\}+\left\{-(D-5)G\right$.
$\left.+(\eta_{7}+\eta_{4}){\partial\over\partial\eta_{4}}G+(\eta_{7}-\eta_{3}+\eta_{5}){\partial\over\partial\eta_{5}}G+2\eta_{7}{\partial\over\partial\eta_{7}}G\right$.
$\left.+(\eta_{7}+\eta_{6}){\partial\over\partial\eta_{6}}G-\eta_{9}{\partial\over\partial\eta_{9}}G\right\}+\left\{{t\over 2s_{0}}\eta_{9}G\right\}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$(43)
$\left\{2{\partial\over\partial\eta_{9}}{\partial\over\partial\eta_{2}}G\right\}=\left\{-(D-5)G+(\eta_{5}+\eta_{1}-\eta_{4}){\partial\over\partial\eta_{1}}G\right$.
$\left.+(\eta_{5}+\eta_{2}-\eta_{4}){\partial\over\partial\eta_{2}}G+(\eta_{5}+\eta_{3}-\eta_{6}){\partial\over\partial\eta_{3}}G\right$.
$\left.+2\eta_{5}{\partial\over\partial\eta_{5}}G-\eta_{8}{\partial\over\partial\eta_{8}}G\right\}+\left\{{s\over 2s_{0}}\eta_{8}G\right\}$
$+\left\{{1\over 2}(\eta_{7}-\eta_{6})\eta_{8}G\right\}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$(44)
$\left\{2{\partial\over\partial\eta_{8}}{\partial\over\partial\eta_{7}}G\right\}=\left\{{s\over s_{0}}{\partial\over\partial\eta_{7}}G\right\}+\left\{-(D-5)G+2\eta_{5}{\partial\over\partial\eta_{5}}G\right$.
$\left.+(\eta_{5}-\eta_{1}+\eta_{4}){\partial\over\partial\eta_{4}}G+(\eta_{5}-\eta_{3}+\eta_{6}){\partial\over\partial\eta_{6}}G\right$.
$\left.+(\eta_{5}-\eta_{3}+\eta_{7}){\partial\over\partial\eta_{7}}G-\eta_{9}{\partial\over\partial\eta_{9}}G\right\}$
$+\left\{{1\over 2}(\eta_{2}-\eta_{1})\eta_{9}G\right\}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$(45)
$2{\partial\over\partial\eta_{8}}{\partial\over\partial\eta_{2}}G=\left\{{s+t\over s_{0}}{\partial\over\partial\eta_{2}}G+{s\over s_{0}}{\partial\over\partial\eta_{3}}G\right$.
$\left.+{4(D-4)s_{0}\over s}{\partial\over\partial\eta_{8}}G+{4s_{0}\over s}\eta_{8}{\partial^{2}\over\partial\eta_{8}^{2}}G\right$.
$\left.-{2s_{0}\over s}(\eta_{1}+\eta_{3}-\eta_{4}+2\eta_{5}-\eta_{6}){\partial\over\partial\eta_{8}}{\partial\over\partial\eta_{5}}G\right$.
$\left.-{2s_{0}\over s}(\eta_{1}+\eta_{3}){\partial\over\partial\eta_{8}}{\partial\over\partial\eta_{3}}G\right$.
$\left.-{2s_{0}\over s}(\eta_{1}+2\eta_{2}+\eta_{3}){\partial\over\partial\eta_{8}}{\partial\over\partial\eta_{2}}G\right$.
$\left.-{2s_{0}\over s}(3\eta_{1}+\eta_{3}){\partial\over\partial\eta_{8}}{\partial\over\partial\eta_{1}}G\right\}+\left\{(D-6)G\right$.
$\left.-(\eta_{4}+\eta_{7}){\partial\over\partial\eta_{4}}G+(\eta_{3}-\eta_{5}-\eta_{6}){\partial\over\partial\eta_{5}}G\right$.
$\left.-(\eta_{6}+\eta_{7}){\partial\over\partial\eta_{6}}G-2\eta_{7}{\partial\over\partial\eta_{7}}G\right$.
$\left.-\eta_{8}{\partial\over\partial\eta_{8}}G+\eta_{9}{\partial\over\partial\eta_{9}}G\right\}+\left\{-{t\over 2s_{0}}\eta_{9}G\right\}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$(46)
$2{\partial\over\partial\eta_{9}}{\partial\over\partial\eta_{7}}=\left\{-{s\over s_{0}}{\partial\over\partial\eta_{6}}G+{t\over s_{0}}{\partial\over\partial\eta_{7}}G+{4s_{0}\over s}\eta_{9}{\partial^{2}\over\partial\eta_{9}^{2}}G\right$.
$\left.+{4(D-4)s_{0}\over s}{\partial\over\partial\eta_{9}}G-{2s_{0}\over s}(3\eta_{4}+\eta_{6}){\partial\over\partial\eta_{9}}{\partial\over\partial\eta_{4}}G\right$.
$\left.+{2s_{0}\over s}(\eta_{1}+\eta_{3}-\eta_{4}-2\eta_{5}-\eta_{6}){\partial\over\partial\eta_{9}}{\partial\over\partial\eta_{5}}G\right$.
$\left.-{2s_{0}\over s}(\eta_{4}+3\eta_{6}){\partial\over\partial\eta_{9}}{\partial\over\partial\eta_{6}}G\right$.
$\left.-{2s_{0}\over s}(\eta_{4}+\eta_{6}+2\eta_{7}){\partial\over\partial\eta_{9}}{\partial\over\partial\eta_{7}}G\right\}$
$+\left\{(D-4)G-(\eta_{1}+\eta_{2}){\partial\over\partial\eta_{1}}G-2\eta_{2}{\partial\over\partial\eta_{2}}G+\eta_{9}{\partial\over\partial\eta_{9}}G\right$.
$\left.-(\eta_{2}+\eta_{3}){\partial\over\partial\eta_{3}}G-(\eta_{1}-\eta_{4}+\eta_{5}){\partial\over\partial\eta_{5}}G+\eta_{8}{\partial\over\partial\eta_{8}}G\right\}$
$+\left\{-{(s+t)\over 2s_{0}}\eta_{8}G\right\}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$(47)
where for simplicity we have neglected the contributions from subsectors.
Now we construct new equations according to step (B) (or more directly, just using the integrability conditions to keep equations up to degree two). After simplification using known reduction rules, when applying to the set ${\cal U}_{2}=(0,0,0,0,0,0,0,n_{8},n_{9})$, the first group, i.e., ${\partial\over\partial\eta_{9}}\left({\partial\over\partial\eta_{8}}{\partial\over\partial\eta_{i}}\right)-{\partial\over\partial\eta_{8}}\left({\partial\over\partial\eta_{9}}{\partial\over\partial\eta_{i}}\right)=0,i=2,5,7$ gives
$0$ $=$ $\left\{{2(D-4)s_{0}\over s}{\partial\over\partial\eta_{8}}{\partial\over\partial\eta_{9}}G+{2s_{0}\over s}\eta_{8}{\partial^{2}\over\partial\eta_{8}^{2}}{\partial\over\partial\eta_{9}}G\right\}$ (48)
$\left\{+{(D-4)\over 2}{\partial\over\partial\eta_{8}}G+{1\over 2}\eta_{8}{\partial^{2}\over\partial\eta_{8}^{2}}G+{1\over 2}\eta_{9}{\partial^{2}\over\partial\eta_{9}^{2}}G\right$.
$\left.-\eta_{8}{\partial\over\partial\eta_{8}}{\partial\over\partial\eta_{9}}G\right\}+R_{1}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$
$0$ $=$ $\left\{+{(D-4)\over 2}{\partial\over\partial\eta_{8}}G+{1\over 2}\eta_{8}{\partial^{2}\over\partial\eta_{8}^{2}}G-{(D-4)\over 2}{\partial\over\partial\eta_{9}}G\right$. (49)
$\left.-{1\over 2}\eta_{9}{\partial^{2}\over\partial\eta_{9}^{2}}G\right\}+R_{2}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$
$0$ $=$ $\left\{{-2(D-4)s_{0}\over s}{\partial\over\partial\eta_{8}}{\partial\over\partial\eta_{9}}G-{2s_{0}\over s}\eta_{9}{\partial^{2}\over\partial\eta_{8}}{\partial^{2}\over\partial\eta_{9}^{2}}G\right\}$ (50)
$\left\{-{(D-4)}{\partial\over\partial\eta_{8}}G-{1\over 2}\eta_{8}{\partial^{2}\over\partial\eta_{8}^{2}}G\right$.
$\left.+{(D-4)\over 2}{\partial\over\partial\eta_{9}}G+{1\over 2}\eta_{9}{\partial^{2}\over\partial\eta_{9}^{2}}G\right$.
$\left.+\left(-\eta_{9}-{4(D-4)s_{0}^{2}\over s^{2}}(\eta_{4}+\eta_{6})\right){\partial\over\partial\eta_{8}}{\partial\over\partial\eta_{9}}G\right$.
$\left.-{4s_{0}^{2}\over s^{2}}(\eta_{4}+\eta_{6})\eta_{9}{\partial\over\partial\eta_{8}}{\partial^{2}\over\partial\eta_{9}^{2}}G\right\}+R_{3}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$
where $R_{i}$ are terms with degree zero or lower. When writing these three equations into the recurrence form, we have
$-R_{1}$ $=$ ${2s_{0}\over s}(n_{8}+1)(n_{9}+1)(D-4+n_{8})F_{\vec{n}+\vec{e}_{8}+\vec{e}_{9}}$ (51)
$+{1\over 2}(n_{8}+1)(D-4+n_{8})F_{\vec{n}+\vec{e}_{8}}$
$+{1\over 2}(n_{9}+1)(n_{9}-2n_{8})F_{\vec{n}+\vec{e}_{9}}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$
$-R_{2}$ $=$ ${1\over 2}(n_{8}+1)(D-4+n_{8})F_{\vec{n}+\vec{e}_{8}}$ (52)
$-{1\over 2}(n_{9}+1)(D-4+n_{9})F_{\vec{n}+\vec{e}_{9}}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$
$R_{3}$ $=$ ${2s_{0}\over s}(n_{8}+1)(n_{9}+1)(D-4+n_{9})F_{\vec{n}+\vec{e}_{8}+\vec{e}_{9}}$ (53)
$+{1\over 2}(n_{8}+1)(2(D-4)+n_{8}+2n_{9})F_{\vec{n}+\vec{e}_{8}}$
$-{1\over 2}(n_{9}+1)(D-4+n_{9})F_{\vec{n}+\vec{e}_{9}}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$
We can solve $F_{\vec{n}+\vec{e}_{8}+\vec{e}_{9}},F_{\vec{n}+\vec{e}_{8}},F_{\vec{n}+\vec{e}_{9}}$ as long as above three relations are linear independent. This can be checked since the coefficient determinant is not zero for any values of $n_{8},n_{9}$.
Equations from the second and third groups are
$0$ $=$ $\left\{{t\over 2s_{0}}{\partial\over\partial\eta_{2}}{\partial\over\partial\eta_{5}}G\right\}+\left\{+{(D-5)\over 2}{\partial\over\partial\eta_{2}}G-\eta_{7}{\partial\over\partial\eta_{7}}{\partial\over\partial\eta_{2}}G\right$. (54)
$\left.+(D-{11\over 2}){\partial\over\partial\eta_{5}}G-\eta_{7}{\partial\over\partial\eta_{7}}{\partial\over\partial\eta_{5}}G-\eta_{5}{\partial^{2}\over\partial\eta_{5}^{2}}G\right$.
$\left.-{t(\eta_{1}+2\eta_{2}+\eta_{3})+s(\eta_{2}+\eta_{5}+\eta_{7})\over 2s}{\partial\over\partial\eta_{2}}{\partial\over\partial\eta_{5}}G\right\}$
$+R_{4}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$
$0$ $=$ $\left\{{t\over 2s_{0}}{\partial\over\partial\eta_{2}}{\partial\over\partial\eta_{7}}G\right\}+\left\{+{(D-5)\over 2}{\partial\over\partial\eta_{2}}G-\eta_{5}{\partial\over\partial\eta_{5}}{\partial\over\partial\eta_{2}}G\right$. (55)
$\left.+(D-{11\over 2}){\partial\over\partial\eta_{7}}G-\eta_{5}{\partial\over\partial\eta_{7}}{\partial\over\partial\eta_{5}}G-\eta_{7}{\partial^{2}\over\partial\eta_{7}^{2}}G\right$.
$\left.-{1\over 2}(\eta_{2}+\eta_{5}+\eta_{7}){\partial\over\partial\eta_{2}}{\partial\over\partial\eta_{7}}G\right\}+R_{5}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$
$0$ $=$ $\left\{+{(D-6)\over 2}{\partial\over\partial\eta_{5}}G-\eta_{5}{\partial^{2}\over\partial\eta_{5}^{2}}G-{(D-6)\over 2}{\partial\over\partial\eta_{7}}G\right$. (56)
$\left.+\eta_{7}{\partial^{2}\over\partial\eta_{7}^{2}}G\right\}+R_{6}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$
$0$ $=$ $\left\{+{(D-6)\over 2}{\partial\over\partial\eta_{2}}G-\eta_{2}{\partial^{2}\over\partial\eta_{2}^{2}}G-{(D-6)\over 2}{\partial\over\partial\eta_{5}}G\right$. (57)
$\left.+\eta_{5}{\partial^{2}\over\partial\eta_{5}^{2}}G\right\}+R_{7}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$
$0$ $=$ $\left\{-{t\over 2s_{0}}{\partial\over\partial\eta_{2}}{\partial\over\partial\eta_{7}}G\right\}+\left\{({11\over 2}-D){\partial\over\partial\eta_{2}}G+\eta_{2}{\partial^{2}\over\partial\eta_{2}^{2}}G\right$. (58)
$\left.-{(D-5)\over 2}{\partial\over\partial\eta_{7}}G+\eta_{5}{\partial\over\partial\eta_{5}}{\partial\over\partial\eta_{7}}G\right$.
$\left.+{1\over 2}(\eta_{2}+\eta_{5}+\eta_{7}){\partial\over\partial\eta_{2}}{\partial\over\partial\eta_{7}}G+\eta_{5}{\partial\over\partial\eta_{2}}{\partial\over\partial\eta_{5}}G\right\}+R_{8}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$
$0$ $=$ $\left\{-{t\over 2s_{0}}{\partial\over\partial\eta_{5}}{\partial\over\partial\eta_{7}}G\right\}+\left\{-(D-{11\over 2}){\partial\over\partial\eta_{5}}G+\eta_{5}{\partial^{2}\over\partial\eta_{5}^{2}}G\right$. (59)
$\left.+\eta_{2}{\partial\over\partial\eta_{2}}{\partial\over\partial\eta_{5}}G-{(D-5)\over 2}{\partial\over\partial\eta_{7}}G+\eta_{2}{\partial\over\partial\eta_{2}}{\partial\over\partial\eta_{7}}G\right$.
$\left.+{1\over 2}(\eta_{2}+\eta_{5}+\eta_{7}){\partial\over\partial\eta_{5}}{\partial\over\partial\eta_{7}}G\right\}+R_{9}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$
Using these $6$ relations, first we can solve three degree two operators ${\partial\over\partial\eta_{2}}{\partial\over\partial\eta_{5}}$, ${\partial\over\partial\eta_{2}}{\partial\over\partial\eta_{7}}$ and ${\partial\over\partial\eta_{5}}{\partial\over\partial\eta_{7}}$. Secondly, we have three degree one relations, which contain three degree one operators with index (0,1,0,0,0,0,0,0,0)(0,1,0,0,0,0,0,0,0),$(0,0,0,0,1,0,0,0,0)$, and $(0,0,0,0,0,0,1,0,0)$. However, they are not independent, as can be seen in the sum of (55), (56), (57), and (58), the degree two and degree one parts have been completely canceled.
Now we can use (56) to solve, for example
${(D-6)\over 2}{\partial\over\partial\eta_{7}}G-\eta_{7}{\partial^{2}\over\partial\eta_{7}^{2}}G$ (60)
$=$ ${(D-6)\over 2}{\partial\over\partial\eta_{5}}G-\eta_{5}{\partial^{2}\over\partial\eta_{5}^{2}}G+R_{6}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$
and
${(D-6)\over 2}{\partial\over\partial\eta_{2}}G-\eta_{2}{\partial^{2}\over\partial\eta_{2}^{2}}G$ (61)
$=$ ${(D-6)\over 2}{\partial\over\partial\eta_{5}}G-\eta_{5}{\partial^{2}\over\partial\eta_{5}^{2}}G-R_{7}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$
Since the left-hand side is the degree one operators, using them, we can reduce all $n_{2},n_{7}$ to zero. After iterating using these two, the unreduced lattice points are just $(0,0,0,0,n_{5},0,0,0,0)$, which contains infinity number of points.
To further reduce points $(0,0,0,0,n_{5},0,0,0,0)$, we need to produce a new relation using newly found reduction rules. Acting $\partial\over\partial\eta_{5}$} at the both sides of (61) we get
${(D-8)\over 2}{\partial^{2}\over\partial\eta_{5}^{2}}G-\eta_{5}{\partial^{3}\over\partial\eta_{5}^{3}}G$ (62)
$=$ ${(D-6)\over 2}{\partial\over\partial\eta_{5}}{\partial\over\partial\eta_{2}}G-\eta_{2}{\partial^{2}\over\partial\eta_{2}^{2}}{\partial\over\partial\eta_{5}}G+{\partial\over\partial\eta_{5}}R_{7}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$
Using (55), the first two terms at the right-hand side can be further reduced, and we get a perfect degree two reduction rule for $(0,0,0,0,n_{5},0,0,0,0)$. Furthermore, since the left hand side is $F_{\vec{n}+2\vec{e}_{5}}$, we can only reduce $n_{5}\geq 2$. Now the only points that can not be reduced are $n_{5}=0,1$ and we have found two master integrals.
Appendix C The top sector of Two-loop massless non-planer
Using the $10$ fundamental IBP relations, we can solve $2$ degree one operators
${s\over s_{0}}{\partial\over\partial\eta_{3}}G=\left\{-2(D-5)G-\eta_{9}{\partial\over\partial\eta_{9}}G-\eta_{8}{\partial\over\partial\eta_{8}}G\right$.
$\left.+2\eta_{7}{\partial\over\partial\eta_{7}}G+(\eta_{6}+\eta_{7}){\partial\over\partial\eta_{6}}G+(\eta_{4}+\eta_{5}){\partial\over\partial\eta_{5}}G+2\eta_{4}{\partial\over\partial\eta_{4}}G\right$.
$\left.+(\eta_{1}+\eta_{3}){\partial\over\partial\eta_{3}}G+(\eta_{1}+\eta_{2}){\partial\over\partial\eta_{2}}G+2\eta_{1}{\partial\over\partial\eta_{1}}G\right\}\penalty 10000\ \penalty 10000\ \penalty 10000\$(63)
${s\over s_{0}}{\partial\over\partial\eta_{1}}G=\left\{-2(D-5)G-\eta_{9}{\partial\over\partial\eta_{9}}G-\eta_{8}{\partial\over\partial\eta_{8}}G\right$.
$\left.+(\eta_{6}+\eta_{7}){\partial\over\partial\eta_{7}}G+2\eta_{6}{\partial\over\partial\eta_{6}}G+2\eta_{5}{\partial\over\partial\eta_{5}}G+(\eta_{4}+\eta_{5}){\partial\over\partial\eta_{4}}G\right$.
$\left.+2\eta_{3}{\partial\over\partial\eta_{3}}G+(\eta_{2}+\eta_{3}){\partial\over\partial\eta_{2}}G+(\eta_{1}+\eta_{3}){\partial\over\partial\eta_{1}}G\right\}$
$+\left\{{-s\eta_{8}-t\eta_{9}\over 2s_{0}}G\right\}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$(64)
and $6$ degree two operators
$2{\partial\over\partial\eta_{6}}{\partial\over\partial\eta_{8}}G=\left\{{-s\over s_{0}}{\partial\over\partial\eta_{6}}G\right\}+\left\{(D-5)G+\eta_{9}{\partial\over\partial\eta_{9}}G\right$.
$\left.+(\eta_{1}-\eta_{4}-\eta_{7}){\partial\over\partial\eta_{7}}G+(\eta_{3}-\eta_{4}-\eta_{6}){\partial\over\partial\eta_{6}}G\right$.
$\left.+(-\eta_{4}-\eta_{5}){\partial\over\partial\eta_{5}}G-2\eta_{4}{\partial\over\partial\eta_{4}}G\right\}\penalty 10000\ \penalty 10000\ \penalty 10000\$(65)
$2{\partial\over\partial\eta_{7}}{\partial\over\partial\eta_{8}}G=\left\{(D-5)G-\eta_{9}{\partial\over\partial\eta_{9}}G+2\eta_{5}{\partial\over\partial\eta_{5}}G\right$.
$\left.+(-\eta_{1}+\eta_{5}+\eta_{7}){\partial\over\partial\eta_{7}}G+(-\eta_{3}+\eta_{5}+\eta_{6}){\partial\over\partial\eta_{6}}G\right$.
$\left.+(\eta_{4}+\eta_{5}){\partial\over\partial\eta_{4}}G\right\}+\left\{-{t\over 2s_{0}}\eta_{9}G\right\}\penalty 10000\ \penalty 10000\ \penalty 10000\$(66)
$2{\partial\over\partial\eta_{5}}{\partial\over\partial\eta_{8}}G=\left\{(D-5)G-\eta_{9}{\partial\over\partial\eta_{9}}G+(2\eta_{7}){\partial\over\partial\eta_{7}}G\right$.
$\left.+(\eta_{6}+\eta_{7}){\partial\over\partial\eta_{6}}G+(-\eta_{1}+\eta_{5}+\eta_{7}){\partial\over\partial\eta_{5}}G\right$.
$\left.+(-\eta_{1}+\eta_{4}+\eta_{7}){\partial\over\partial\eta_{4}}G\right\}+\left\{{\eta_{1}-\eta_{2}\over 2}\eta_{9}G\right\}\penalty 10000\ \penalty 10000\ \penalty 10000\$(67)
$2{\partial\over\partial\eta_{4}}{\partial\over\partial\eta_{8}}G=\left\{{-s\over s_{0}}{\partial\over\partial\eta_{4}}G\right\}+\left\{(D-5)G+\eta_{9}{\partial\over\partial\eta_{9}}G\right$.
$\left.+(-\eta_{6}-\eta_{7}){\partial\over\partial\eta_{7}}G+(-2\eta_{6}){\partial\over\partial\eta_{6}}G+(-\eta_{4}+\eta_{6}){\partial\over\partial\eta_{4}}G\right$.
$\left.+(\eta_{3}-\eta_{5}-\eta_{6}){\partial\over\partial\eta_{5}}G\right\}+\left\{{(s+t)\over 2s_{0}}\eta_{9}G\right\}$
$+\left\{{-\eta_{1}+\eta_{2}\over 2}\eta_{9}G\right\}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$(68)
$2{\partial\over\partial\eta_{2}}{\partial\over\partial\eta_{8}}G=\left\{{4(2D-9)s_{0}\over s}{\partial\over\partial\eta_{8}}G+{4s_{0}\eta_{9}\over s}{\partial\over\partial\eta_{8}}{\partial\over\partial\eta_{9}}G\right$.
$\left.+{4s_{0}\eta_{8}\over s}{\partial^{2}\over\partial\eta_{8}^{2}}G-{2s_{0}(\eta_{6}+3\eta_{7})\over s}{\partial\over\partial\eta_{7}}{\partial\over\partial\eta_{8}}G\right$.
$-{2s_{0}(3\eta_{6}+\eta_{7})\over s}{\partial\over\partial\eta_{6}}{\partial\over\partial\eta_{8}}G-{2s_{0}(3\eta_{5}+\eta_{4})\over s}{\partial\over\partial\eta_{5}}{\partial\over\partial\eta_{8}}G$
$-{2s_{0}(3\eta_{4}+\eta_{5})\over s}{\partial\over\partial\eta_{4}}{\partial\over\partial\eta_{8}}-{2s_{0}(3\eta_{3}+\eta_{1})\over s}{\partial\over\partial\eta_{3}}{\partial\over\partial\eta_{8}}G$
$+{t\over s_{0}}{\partial\over\partial\eta_{2}}G-{2s_{0}(\eta_{1}+2\eta_{2}+\eta_{3})\over s}{\partial\over\partial\eta_{2}}{\partial\over\partial\eta_{8}}G$
$\left.-{s\over s_{0}}{\partial\over\partial\eta_{1}}G-{2s_{0}(3\eta_{1}+\eta_{3})\over s}{\partial\over\partial\eta_{1}}{\partial\over\partial\eta_{8}}G\right\}$
$+\left\{G+{t\over s}\eta_{9}{\partial\over\partial\eta_{8}}G+(\eta_{6}-\eta_{7}){\partial\over\partial\eta_{7}}G+(\eta_{6}-\eta_{7}){\partial\over\partial\eta_{6}}G\right$.
$\left.+(\eta_{3}-\eta_{1}){\partial\over\partial\eta_{3}}G+(-\eta_{1}+\eta_{3}){\partial\over\partial\eta_{2}}G\right$.
$\left.+(-\eta_{1}+\eta_{3}){\partial\over\partial\eta_{1}}G\right\}+\left\{{-s\over 2s_{0}}\eta_{8}G\right\}$
$+\left\{{-\eta_{1}+\eta_{2}\over 2}\eta_{9}G\right\}\penalty 10000\ \penalty 10000\ \penalty 10000\$(69)
$2{\partial\over\partial\eta_{2}}{\partial\over\partial\eta_{9}}G=\left\{{2(2D-9)s_{0}\over s}{\partial\over\partial\eta_{8}}G+{2s_{0}\eta_{9}\over s}{\partial\over\partial\eta_{8}}{\partial\over\partial\eta_{9}}G\right$.
$\left.+{2s_{0}\eta_{8}\over s}{\partial^{2}\over\partial\eta_{8}^{2}}G-{4s_{0}\eta_{7}\over s}{\partial\over\partial\eta_{7}}{\partial\over\partial\eta_{8}}G\right$.
$-{2s_{0}(\eta_{6}+\eta_{7})\over s}{\partial\over\partial\eta_{6}}{\partial\over\partial\eta_{8}}G-{2s_{0}(\eta_{5}+\eta_{4})\over s}{\partial\over\partial\eta_{5}}{\partial\over\partial\eta_{8}}G$
$-{4s_{0}\eta_{4}\over s}{\partial\over\partial\eta_{4}}{\partial\over\partial\eta_{8}}G-{2s_{0}(\eta_{1}+\eta_{3})\over s}{\partial\over\partial\eta_{3}}{\partial\over\partial\eta_{8}}G+{-s\over s_{0}}{\partial\over\partial\eta_{3}}G$
$\left.-{2s_{0}(\eta_{1}+\eta_{2})\over s}{\partial\over\partial\eta_{2}}{\partial\over\partial\eta_{8}}G-{4s_{0}\eta_{1}\over s}{\partial\over\partial\eta_{1}}{\partial\over\partial\eta_{8}}G\right\}$
$+\left\{-(D-5)G-\eta_{9}{\partial\over\partial\eta_{9}}G+(\eta_{4}+\eta_{5}){\partial\over\partial\eta_{5}}G+2\eta_{4}{\partial\over\partial\eta_{4}}G\right$.
$\left.+(\eta_{3}+\eta_{5}-\eta_{6}){\partial\over\partial\eta_{3}}G+(\eta_{1}+\eta_{4}-\eta_{7}){\partial\over\partial\eta_{2}}G\right$.
$\left.+(\eta_{1}+\eta_{4}-\eta_{7}){\partial\over\partial\eta_{1}}G\right\}+\left\{{\eta_{4}-\eta_{5}\over 2}\eta_{8}G\right\}\penalty 10000\ \penalty 10000\ \penalty 10000\$(70)
Adding two new integrability conditions, i.e., ${\partial\over\partial\eta_{2}}\left({\partial\over\partial\eta_{4}}+{\partial\over\partial\eta_{5}}\right){\partial\over\partial\eta_{9}}-\left({\partial\over\partial\eta_{4}}+{\partial\over\partial\eta_{5}}\right){\partial\over\partial\eta_{2}}{\partial\over\partial\eta_{9}}=0$ and ${\partial\over\partial\eta_{2}}\left({\partial\over\partial\eta_{6}}+{\partial\over\partial\eta_{7}}\right){\partial\over\partial\eta_{9}}-\left({\partial\over\partial\eta_{6}}+{\partial\over\partial\eta_{7}}\right){\partial\over\partial\eta_{2}}{\partial\over\partial\eta_{9}}=0$, we can solve following four degree two operators
$-{\partial\over\partial\eta_{4}}{\partial\over\partial\eta_{9}}G=\left\{{(9-2D)s_{0}\over s}{\partial\over\partial\eta_{8}}G-{s_{0}\over s}\eta_{9}{\partial\over\partial\eta_{8}}{\partial\over\partial\eta_{9}}G\right$.
$\left.-{s_{0}\over s}\eta_{8}{\partial^{2}\over\partial\eta_{8}^{2}}G+{s_{0}\over s}(\eta_{6}-\eta_{7}){\partial\over\partial\eta_{7}}{\partial\over\partial\eta_{9}}G+{2s_{0}\over s}\eta_{7}{\partial\over\partial\eta_{7}}{\partial\over\partial\eta_{8}}G\right$.
$+{s_{0}\over s}(\eta_{6}-\eta_{7}){\partial\over\partial\eta_{6}}{\partial\over\partial\eta_{9}}G+{s_{0}\over s}(\eta_{6}+\eta_{7}){\partial\over\partial\eta_{6}}{\partial\over\partial\eta_{8}}G$
$+{s_{0}\over s}(\eta_{1}-\eta_{3}+\eta_{6}-\eta_{7}){\partial\over\partial\eta_{5}}{\partial\over\partial\eta_{9}}G+{s_{0}\over s}(\eta_{4}+\eta_{5}){\partial\over\partial\eta_{5}}{\partial\over\partial\eta_{8}}G$
$+{s_{0}\over s}(\eta_{1}-\eta_{3}+\eta_{6}-\eta_{7}){\partial\over\partial\eta_{4}}{\partial\over\partial\eta_{9}}G+{2s_{0}\over s}\eta_{4}{\partial\over\partial\eta_{4}}{\partial\over\partial\eta_{8}}G$
$\left.+{(\eta_{2}+\eta_{3})s_{0}\over s}{\partial\over\partial\eta_{3}}{\partial\over\partial\eta_{8}}G+{2s_{0}\over s}\eta_{2}{\partial\over\partial\eta_{2}}{\partial\over\partial\eta_{8}}G\right$.
$\left.+{s_{0}(\eta_{1}+\eta_{2})\over s}{\partial\over\partial\eta_{1}}{\partial\over\partial\eta_{8}}G\right\}+\left\{{-2s_{0}+(D-5)t\over 2s}G-{\eta_{9}\over 2}{\partial\over\partial\eta_{9}}G\right$.
$\left.+{t\eta_{8}\over 2s}{\partial\over\partial\eta_{8}}G+{-t\eta_{7}\over s}{\partial\over\partial\eta_{7}}G+{-t(\eta_{6}+\eta_{7})\over 2s}{\partial\over\partial\eta_{6}}G\right$.
$\left.+{t(\eta_{1}-\eta_{5}-\eta_{7})\over 2s}{\partial\over\partial\eta_{5}}G+{t(\eta_{1}-\eta_{4}-\eta_{7})\over 2s}{\partial\over\partial\eta_{4}}G\right\}$
$+\left\{+{t(-\eta_{1}+\eta_{2})\eta_{9}\over 4s}G\right\}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$(71)
$-{\partial\over\partial\eta_{5}}{\partial\over\partial\eta_{9}}G=\left\{{-(9-2D)s_{0}\over s}{\partial\over\partial\eta_{8}}G+{s_{0}\over s}\eta_{9}{\partial\over\partial\eta_{8}}{\partial\over\partial\eta_{9}}G\right$.
$\left.+{s_{0}\over s}\eta_{8}{\partial^{2}\over\partial\eta_{8}^{2}}G-{s_{0}\over s}(\eta_{6}-\eta_{7}){\partial\over\partial\eta_{7}}{\partial\over\partial\eta_{9}}G-{2s_{0}\over s}\eta_{7}{\partial\over\partial\eta_{7}}{\partial\over\partial\eta_{8}}G\right$.
$-{s_{0}\over s}(\eta_{6}-\eta_{7}){\partial\over\partial\eta_{6}}{\partial\over\partial\eta_{9}}G-{s_{0}\over s}(\eta_{6}+\eta_{7}){\partial\over\partial\eta_{6}}{\partial\over\partial\eta_{8}}G$
$+{t\over 2s_{0}}{\partial\over\partial\eta_{5}}G-{s_{0}\over s}(\eta_{1}-\eta_{3}+\eta_{6}-\eta_{7}){\partial\over\partial\eta_{5}}{\partial\over\partial\eta_{9}}G$
$-{s_{0}\over s}(\eta_{4}+\eta_{5}){\partial\over\partial\eta_{5}}{\partial\over\partial\eta_{8}}G-{2s_{0}\over s}\eta_{4}{\partial\over\partial\eta_{4}}{\partial\over\partial\eta_{8}}G$
$-{s_{0}(\eta_{1}-\eta_{3}+\eta_{6}-\eta_{7})\over s}{\partial\over\partial\eta_{4}}{\partial\over\partial\eta_{9}}G+{-s\over 2s_{0}}{\partial\over\partial\eta_{3}}G$
$\left.-{(\eta_{2}+\eta_{3})s_{0}\over s}{\partial\over\partial\eta_{3}}{\partial\over\partial\eta_{8}}G-{2s_{0}\over s}\eta_{2}{\partial\over\partial\eta_{2}}{\partial\over\partial\eta_{8}}G\right$.
$\left.-{s_{0}(\eta_{1}+\eta_{2})\over s}{\partial\over\partial\eta_{1}}{\partial\over\partial\eta_{8}}G\right\}+\left\{{s+(5-D)t\over 2s}G+{\eta_{9}\over 2}{\partial\over\partial\eta_{9}}G\right$.
$\left.-{t\eta_{8}\over 2s}{\partial\over\partial\eta_{8}}G+{t\eta_{7}\over s}{\partial\over\partial\eta_{7}}G+{t(\eta_{6}+\eta_{7})\over 2s}{\partial\over\partial\eta_{6}}G\right$.
$\left.-{t(\eta_{1}-\eta_{5}-\eta_{7})\over 2s}{\partial\over\partial\eta_{5}}G+{-t(\eta_{1}-\eta_{4}-\eta_{7})\over 2s}{\partial\over\partial\eta_{4}}G\right$.
$\left.+{(\eta_{1}-\eta_{2})\over 2s}{\partial\over\partial\eta_{3}}G+{(\eta_{1}-\eta_{2})\over 2s}{\partial\over\partial\eta_{2}}G+{(\eta_{1}-\eta_{2})\over 2s}{\partial\over\partial\eta_{1}}G\right\}$
$+\left\{-{t\over 4s_{0}}\eta_{8}G\right\}+\left\{{-t(-\eta_{1}+\eta_{2})\eta_{9}\over 4s}G\right\}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$(72)
$-{\partial\over\partial\eta_{6}}{\partial\over\partial\eta_{9}}G=\left\{{(2D-9)s_{0}\over s}{\partial\over\partial\eta_{8}}G+{s_{0}\eta_{9}\over s}{\partial\over\partial\eta_{8}}{\partial\over\partial\eta_{9}}G\right$.
$\left.+{s_{0}\eta_{8}\over s}{\partial^{2}\over\partial\eta_{8}^{2}}G+{s_{0}(\eta_{4}-\eta_{5})\over s}{\partial\over\partial\eta_{7}}{\partial\over\partial\eta_{9}}G\right$.
$+{s_{0}(\eta_{1}-\eta_{2}-2\eta_{7})\over s}\eta_{7}{\partial\over\partial\eta_{7}}{\partial\over\partial\eta_{8}}G+{(s+t)\over 2s_{0}}{\partial\over\partial\eta_{6}}G$
$+{s_{0}(\eta_{4}-\eta_{5})\over s}{\partial\over\partial\eta_{6}}{\partial\over\partial\eta_{9}}G-{s_{0}(-\eta_{1}+\eta_{2}+\eta_{6}+\eta_{7})\over s}{\partial\over\partial\eta_{6}}{\partial\over\partial\eta_{8}}G$
$+{s_{0}(\eta_{4}-\eta_{5})\over s}{\partial\over\partial\eta_{5}}{\partial\over\partial\eta_{9}}G-{s_{0}(\eta_{4}+\eta_{5})\over s}{\partial\over\partial\eta_{5}}{\partial\over\partial\eta_{8}}G$
$+{s_{0}(\eta_{4}-\eta_{5})\over s}{\partial\over\partial\eta_{4}}{\partial\over\partial\eta_{9}}G-{2s_{0}\eta_{4}\over s}{\partial\over\partial\eta_{4}}{\partial\over\partial\eta_{8}}G$
$-{(\eta_{2}+\eta_{3})s_{0}\over s}{\partial\over\partial\eta_{3}}{\partial\over\partial\eta_{8}}G-{2s_{0}\eta_{2}\over s}{\partial\over\partial\eta_{2}}{\partial\over\partial\eta_{8}}G$
$\left.-{s_{0}(\eta_{1}+\eta_{2})\over s}{\partial\over\partial\eta_{1}}{\partial\over\partial\eta_{8}}G\right\}+\left\{{-(D-5)(s+t)\over 2s}G-{\eta_{9}\over 2}{\partial\over\partial\eta_{9}}G\right$.
$\left.+{-t\eta_{8}\over 2s}{\partial\over\partial\eta_{8}}G+{-(s+t)(\eta_{1}-\eta_{4}-\eta_{7})\over 2s}{\partial\over\partial\eta_{7}}G\right$.
$\left.+{-(s+t)(\eta_{3}-\eta_{4}-\eta_{6})\over 2s}{\partial\over\partial\eta_{6}}G+{(s+t)(\eta_{4}+\eta_{5})\over 2s}{\partial\over\partial\eta_{5}}G\right$.
$\left.+{(s+t)\eta_{4}\over s}{\partial\over\partial\eta_{4}}G\right\}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$(73)
and
$-{\partial\over\partial\eta_{7}}{\partial\over\partial\eta_{9}}G=\left\{{-(2D-9)s_{0}\over s}{\partial\over\partial\eta_{8}}G-{s_{0}\eta_{9}\over s}{\partial\over\partial\eta_{8}}{\partial\over\partial\eta_{9}}G\right$.
$\left.-{s_{0}\eta_{8}\over s}{\partial^{2}\over\partial\eta_{8}^{2}}G-{s_{0}(\eta_{4}-\eta_{5})\over s}{\partial\over\partial\eta_{7}}{\partial\over\partial\eta_{9}}G\right$.
$-{s_{0}(\eta_{1}-\eta_{2}-2\eta_{7})\over s}\eta_{7}{\partial\over\partial\eta_{7}}{\partial\over\partial\eta_{8}}G-{s_{0}(\eta_{4}-\eta_{5})\over s}{\partial\over\partial\eta_{6}}{\partial\over\partial\eta_{9}}G$
$+{s_{0}(-\eta_{1}+\eta_{2}+\eta_{6}+\eta_{7})\over s}{\partial\over\partial\eta_{6}}{\partial\over\partial\eta_{8}}G-{s_{0}(\eta_{4}-\eta_{5})\over s}{\partial\over\partial\eta_{5}}{\partial\over\partial\eta_{9}}G$
$+{s_{0}(\eta_{4}+\eta_{5})\over s}{\partial\over\partial\eta_{5}}{\partial\over\partial\eta_{8}}G-{s_{0}(\eta_{4}-\eta_{5})\over s}{\partial\over\partial\eta_{4}}{\partial\over\partial\eta_{9}}G$
$+{2s_{0}\eta_{4}\over s}{\partial\over\partial\eta_{4}}{\partial\over\partial\eta_{8}}G+{s\over 2s_{0}}{\partial\over\partial\eta_{3}}G+{(\eta_{2}+\eta_{3})s_{0}\over s}{\partial\over\partial\eta_{3}}{\partial\over\partial\eta_{8}}G$
$\left.+{2s_{0}\eta_{2}\over s}{\partial\over\partial\eta_{2}}{\partial\over\partial\eta_{8}}G+{s_{0}(\eta_{1}+\eta_{2})\over s}{\partial\over\partial\eta_{1}}{\partial\over\partial\eta_{8}}G\right\}$
$+\left\{{(D-5)(s+t)\over 2s}G+{\eta_{9}\over 2}{\partial\over\partial\eta_{9}}G+{t\eta_{8}\over 2s}{\partial\over\partial\eta_{8}}G\right$.
$\left.-{t(-\eta_{1}+\eta_{4}+\eta_{7})+s(-\eta_{2}+\eta_{4}+\eta_{7})\over 2s}{\partial\over\partial\eta_{7}}G\right$.
$\left.+{t(\eta_{3}-\eta_{4}-\eta_{6})-s(\eta_{1}-\eta_{2}-\eta_{3}+\eta_{4}+\eta_{6})\over 2s}{\partial\over\partial\eta_{6}}G\right$.
$\left.-{(s+t)(\eta_{4}+\eta_{5})\over 2s}{\partial\over\partial\eta_{5}}G-{(s+t)\eta_{4}\over s}{\partial\over\partial\eta_{4}}G\right$.
$\left.+{(-\eta_{1}+\eta_{2})\over 2}{\partial\over\partial\eta_{3}}G+{(-\eta_{1}+\eta_{2})\over 2}{\partial\over\partial\eta_{2}}G\right$.
$\left.+{(-\eta_{1}+\eta_{2})\over 2}{\partial\over\partial\eta_{1}}G\right\}+\left\{{t\over 4s_{0}}\eta_{8}G\right\}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$(74)
Using above reduction rules, the sets can not be reduced are
${\cal U}_{1}$ $=$ $(0,n_{2},0,n_{4},n_{5},n_{6},n_{7},0,0),$
${\cal U}_{2}$ $=$ $(0,0,0,0,0,0,0,n_{8},n_{9})$ (75)
Now we consider the new equations coming from integrability condition ${\partial\over\partial\eta_{9}}\left({\partial\over\partial\eta_{i}}{\partial\over\partial\eta_{8}}\right)-{\partial\over\partial\eta_{8}}\left({\partial\over\partial\eta_{i}}{\partial\over\partial\eta_{9}}\right)=0,i=2,4,5,6,7$. After proper combination, they give following three useful equations
$0$ $=$ $\left\{+{2(D-4)s_{0}\over s}{\partial^{2}\over\partial\eta_{8}^{2}}G+{s_{0}\over s}\eta_{8}{\partial^{3}\over\partial\eta_{8}^{3}}G+{s_{0}\over s}\eta_{9}{\partial^{2}\over\partial\eta_{8}^{2}}{\partial\over\partial\eta_{9}}G\right\}$ (76)
$+\left\{{(4-D)\over 2}{\partial\over\partial\eta_{9}}G-{\eta_{9}\over 2}{\partial^{2}\over\partial\eta_{9}^{2}}G\right$.
$\left.+\left((D-4)-{(D-4)t\over 2s}\right){\partial\over\partial\eta_{8}}G+\eta_{9}{\partial\over\partial\eta_{8}}{\partial\over\partial\eta_{9}}G\right$.
$\left.+{4(D-4)(\eta_{1}-2\eta_{2}+\eta_{3})s_{0}^{2}+s(s-t)\eta_{8}\over 2s^{2}}{\partial^{2}\over\partial\eta_{8}^{2}}G\right$.
$\left.+{s_{0}^{2}(\eta_{1}-2\eta_{2}+\eta_{3})\eta_{9}\over s^{2}}{\partial^{2}\over\partial\eta_{8}^{2}}{\partial\over\partial\eta_{9}}G\right$.
$\left.+{s_{0}^{2}(\eta_{1}-2\eta_{2}+\eta_{3})\eta_{8}\over s^{2}}{\partial^{3}\over\partial\eta_{8}^{3}}G\right\}+R_{489}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$
$0$ $=$ $\left\{-{4(D-4)s_{0}\over s}{\partial\over\partial\eta_{8}}{\partial\over\partial\eta_{9}}G-{2s_{0}\over s}\eta_{9}{\partial\over\partial\eta_{8}}{\partial^{2}\over\partial\eta_{9}^{2}}G\right$. (77)
$\left.+{-2s_{0}\over s}\eta_{8}{\partial^{2}\over\partial\eta_{8}^{2}}{\partial\over\partial\eta_{9}}G\right\}+\left\{{(4-D)\over 2}{\partial\over\partial\eta_{9}}G\right$.
$\left.-{(D-4)(s+t)\over 2s}{\partial\over\partial\eta_{8}}G\right$.
$\left.+{-4(D-4)(\eta_{1}-2\eta_{2}+\eta_{3})s_{0}^{2}+s(s\eta_{8}+(s+t)\eta_{9})\over s^{2}}\right$.
$\left.{\partial\over\partial\eta_{8}}{\partial\over\partial\eta_{9}}G+{-2s_{0}^{2}(\eta_{1}-2\eta_{2}+\eta_{3})\eta_{9}\over s^{2}}{\partial\over\partial\eta_{8}}{\partial^{2}\over\partial\eta_{9}^{2}}G\right$.
$\left.+{-2s_{0}^{2}(\eta_{1}-2\eta_{2}+\eta_{3})\eta_{8}\over s^{2}}{\partial^{2}\over\partial\eta_{8}^{2}}{\partial\over\partial\eta_{9}}G\right\}+R_{2489}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$
$0$ $=$ $\left\{(D-4){\partial\over\partial\eta_{9}}G+\eta_{9}{\partial^{2}\over\partial\eta_{9}^{2}}G+{(4-D)\over 2}{\partial\over\partial\eta_{8}}G\right$. (78)
$\left.-\eta_{9}{\partial\over\partial\eta_{8}}{\partial\over\partial\eta_{9}}G\right\}+R_{589}-R_{689}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$
Although the above three equations are enough to finish the reduction for the set ${\cal U}_{2}$, to see it easier, we act $\partial\over\partial\eta_{8}$} on (78), which produces
$0$ $=$ $\left\{(D-4){\partial\over\partial\eta_{8}}{\partial\over\partial\eta_{9}}G+\eta_{9}{\partial\over\partial\eta_{8}}{\partial^{2}\over\partial\eta_{9}^{2}}G+{(4-D)\over 2}{\partial^{2}\over\partial\eta_{8}^{2}}G\right$. (79)
$\left.-\eta_{9}{\partial^{2}\over\partial\eta_{8}^{2}}{\partial\over\partial\eta_{9}}G\right\}+\left\{+{(D-4)t\over 4s_{0}}{\partial\over\partial\eta_{8}}G+{t\eta_{9}\over 2s_{0}}{\partial\over\partial\eta_{8}}{\partial\over\partial\eta_{9}}G\right$.
$\left.-{2(D-4)(\eta_{4}-\eta_{5}-\eta_{6}+\eta_{7})s_{0}\over s}{\partial^{2}\over\partial\eta_{8}^{2}}G\right$.
$\left.+{(-\eta_{4}+\eta_{5}+\eta_{6}-\eta_{7})\eta_{9}s_{0}\over s}{\partial^{2}\over\partial\eta_{8}^{2}}{\partial\over\partial\eta_{9}}G\right$.
$\left.+{(-\eta_{4}+\eta_{5}+\eta_{6}-\eta_{7})\eta_{8}s_{0}\over s}{\partial^{3}\over\partial\eta_{8}^{3}}\right\}+R_{5689}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$
Now we write down the recurrence relations of (77), (76), (78) and (79) for the set $(0,0,0,0,0,0,0,n_{8},n_{9})$. With this constraint, many terms with multiplication factor $\eta_{i},i=1,2,3,4,5,6,7$ will be zero. The four recurrence relations are
$-R_{2489}={s_{0}(n_{9}+1)(n_{8}+1)\over s}(-4(D-4)-2n_{9}-2n_{8})$
$\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ F_{\vec{n}+\vec{e}_{8}+\vec{e}_{9}}+(n_{9}+1)({(4-D)\over 2}+n_{8})F_{\vec{n}+\vec{e}_{9}}$
$\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ +(n_{8}+1){(s+t)\over s}(-{(D-4)\over 2}+n_{9})F_{\vec{n}+\vec{e}_{8}}$
$-R_{489}={s_{0}(n_{8}+2)(n_{8}+1)\over s}(2(D-4)+n_{8}+n_{9})$
$\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ F_{\vec{n}+2\vec{e}_{8}}+{n_{9}+1\over 2}(4-D-n_{9})F_{\vec{n}+\vec{e}_{9}}$
$\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ +(n_{8}+1)({2s-t\over 2s}(D-4)+n_{9}+{s-t\over 2s}n_{8})F_{\vec{n}+\vec{e}_{8}}$
$-R_{589}-R_{689}=(n_{9}+1)(D-4+n_{9})F_{\vec{n}+\vec{e}_{9}}$
$\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ +(n_{8}+1)({(4-D)\over 2}-n_{9})F_{\vec{n}+\vec{e}_{8}}$
$-R_{5689}=(n_{9}+1)(n_{8}+1)(D-4+n_{9})F_{\vec{n}+\vec{e}_{8}+\vec{e}_{9}}$
$\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ +(n_{8}+2)(n_{8}+1)({4-D\over 2}-n_{9})F_{\vec{n}+2\vec{e}_{8}}$
$\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ +(n_{8}+1){t\over 4s_{0}}((D-4)+2n_{9})F_{\vec{n}+\vec{e}_{8}}$ (80)
One can check that determinant of coefficients of above four equations for $\{F_{\vec{n}+\vec{e}_{8}+\vec{e}_{9}},F_{\vec{n}+2\vec{e}_{8}},F_{\vec{n}+\vec{e}_{9}},F_{\vec{n}+\vec{e}_{8}}\}$ is nonzero for general choice of $s,t$. Thus, using four independent relations, we can solve $\{F_{\vec{n}+\vec{e}_{8}+\vec{e}_{9}},F_{\vec{n}+2\vec{e}_{8}},F_{\vec{n}+\vec{e}_{9}},F_{\vec{n}+\vec{e}_{8}}\}$, which give the reduction rules for the set $(0,0,0,0,0,0,0,n_{8},n_{9})$.
Now we consider new equations coming from integrability condition ${\partial\over\partial\eta_{j}}\left({\partial\over\partial\eta_{i}}{\partial\over\partial\eta_{k}}\right)-{\partial\over\partial\eta_{i}}\left({\partial\over\partial\eta_{j}}{\partial\over\partial\eta_{k}}\right)=0,i<j,i,j=2,4,5,6,7;k=8,9$. There are $20$ equations, but some are not independent. Using them, we can solve $10$ degree two operators ${\partial\over\partial\eta_{j}}{\partial\over\partial\eta_{i}}$, $i,j\in\{2,4,5,6,7\},i<j$. We will also have four degree one equations, and using them, we can solve
$\left\{(3-D/2){\partial\over\partial\eta_{6}}G+\eta_{6}{\partial^{2}\over\partial\eta_{6}^{2}}G\right\}$ (81)
$=$ $\left\{(3-D/2){\partial\over\partial\eta_{4}}G+\eta_{4}{\partial^{2}\over\partial\eta_{4}^{2}}G\right\}-R_{4868}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$
$\left\{(3-D/2){\partial\over\partial\eta_{7}}G+\eta_{7}{\partial^{2}\over\partial\eta_{7}^{2}}G\right\}$ (82)
$=$ $\left\{+(3-D/2){\partial\over\partial\eta_{5}}G+\eta_{5}{\partial^{2}\over\partial\eta_{5}^{2}}G\right\}+R_{5878}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$
$\left\{(3-D/2){\partial\over\partial\eta_{2}}G+\eta_{2}{\partial^{2}\over\partial\eta_{2}^{2}}G\right\}$ (83)
$=$ $\left\{+(3-D/2){\partial\over\partial\eta_{5}}G+\eta_{5}{\partial^{2}\over\partial\eta_{5}^{2}}G\right$.
$\left.+(3-D/2){\partial\over\partial\eta_{4}}G+\eta_{4}{\partial^{2}\over\partial\eta_{4}^{2}}G\right\}-R_{29459}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$
$\left\{+(3-D/2){\partial\over\partial\eta_{5}}G+\eta_{5}{\partial^{2}\over\partial\eta_{5}^{2}}G\right\}$ (84)
$=$ ${t+s\over t}\left\{+(3-D/2){\partial\over\partial\eta_{4}}G+\eta_{4}{\partial^{2}\over\partial\eta_{4}^{2}}G\right\}$
$+{s\over t}R_{4979}-R_{5878}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$
Up to now, we see that the points that can not be reduced are $(0,0,0,n_{4},0,0,0,0,0)$. To reduce the $n_{4}$, we can take the $\partial\over\partial\eta_{4}$} at both sides of any degree one relation of the above three, for example, (84). Using the known reduction rule of ${\partial\over\partial\eta_{4}}{\partial\over\partial\eta_{5}}G$ we can reduce the degree on the left-hand side by one, while the right-hand side will have $\partial^{2}\over\partial\eta_{4}^{2}$}, which is a degree two relation for $n_{4}$. Using it, we can reduce $n_{4}$ to $0,1$, thus we get two master integrals.